A Complete Bibliography of the Journal of Graphics Tools: JGT and Journal of Graphics, GPU, and Game

Tools

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 2.12

Title word cross-reference

2 [237, 181, 154, 183]. 3 [82, 168, 79, 78, 114, 291, 4, 181, 154, 312, 295, 109, 99, 215]. 4 [73]. 8 [73]. A [55]. N [48]. T [5]. Z [9].

- [73]. -Buffer [55]. -Buffering [9]. -D [183]. -rooks [48]. -Subdivision [215]. -vertices [5].

2014 [303].

3D [137].

AABB [28]. Accelerated [160, 88].

Accelerating [239, 63, 67]. Accumulation [56]. accuracy [4]. Accurate [269, 40, 63, 268, 95, 99, 8]. Acoustical [274]. Active [195]. Adapting [311]. Adaptive [260, 159, 205, 22, 71, 215, 145]. Adjacency [163, 80]. Advanced [193]. Affine [263]. Algorithm [309, 241, 137, 269, 104, 29, 129, 58, 230, 191, 298, 203, 281, 170, 272, 97, 77, 158, 4, 85, 83, 7]. Algorithms [64, 161]. Alias [202]. Alias-Free [202]. Aligned [217, 142]. Alternative [281]. Ambient [74, 209]. among [117]. Analysis [268]. Animation [254]. Anisotropic [69]. antialiased [13]. Antialiasing [55, 169, 18]. Applications [231]. Approach [292, 176, 210]. Approximate [274, 176, 201, 276, 38]. Approximation [196, 305]. Arbitrary [149]. Arbitrary-Degree [149]. Archive [309]. Arcs [264, 301]. Area [233, 123, 52, 103]. Articulated [143]. Artifacts [263, 249]. Artistic [177]. Artists [295]. As-Rigid-As-Possible [240]. Ashikhmin [119]. Assets [150]. Assisted [159, 24, 116]. Attributes [149]. Authoring [313]. Automatic [152]. Avoiding [2]. Axis [142].

B [302, 294, 40, 63, 211, 238, 157]. B-spline [302, 211, 157, 294, 40, 63]. B-Splines [238]. Backface [32, 266]. Backprojection [221]. bad [11]. Balls [305]. Band [284]. Banishing [11]. Barycentric [167, 297, 98]. Bas [24]. Bas- [24]. Based [160, 241, 194, 60, 274, 176, 304, 43, 130, 58, 201, 213, 135, 93, 169, 238, 141, 51, 23, 311, 154, 1, 87, 145]. Basis [291, 61]. Batchability [312]. Bayer [234]. Between [144, 181, 143, 27, 245]. beyond [121]. Bézier [184, 108, 111]. Bidirectional [191, 215]. Bijective [297]. Bilateral [239]. Bilinear [148]. Billboard [179]. Binary [251]. Binary-Scene [251]. Blendshapes [292]. Blob [96]. Blur [266, 199]. blurry [112]. BMRT [12]. Board [279, 296, 314]. Bodies [304, 143]. Boundaries [211, 183]. Boundary [231]. Bounding [64, 197, 217, 267, 195, 142, 52]. Bounds [108, 310]. Box [217, 198, 142, 158, 82]. Boxes [64, 52]. Branching [243]. BRDF [69, 119]. Breadth [94]. Breadth-First [94]. Broad [248]. Broad-Phase [248]. Browser [242]. BSP [29]. buckling [11]. Buffer [132, 58, 55, 56, 19, 2]. Buffering [43, 9, 276]. Building [291, 61, 57, 252]. Bump [261]. Bumpmap [114].

C [152]. CAB [143]. Calculating [237, 178, 181]. Calculations [93]. Camera [76, 172]. Cases [127]. Casters [230]. Casting [194, 310, 46]. Catmull [211].

Caustics [220]. CCD [293]. Cells [265]. Center [237]. Champagnat [302]. characteristics [13]. Choosing [31]. Chordlength [37]. ChoreoGraphics [313]. ChromaDepth [39]. Cinematography [16]. Circle [298, 99]. Circle-Circle [99]. Circle-Line [99]. Circular [264, 301]. Clark [211]. Classification [170]. clears [2]. Clipping [155, 275, 281, 229, 79]. Close [19]. Closest [117]. Cloud [147]. Clouds [250, 179, 78]. Cluster [224]. Clustered [32]. Coherent [247, 106]. Collection [30]. Collision [60, 176, 47, 248, 143, 174, 28, 50]. Color [236, 226, 272, 31, 1]. Color-to-Gray [226]. Colors [31]. Commodity [122]. Compact [299, 80]. Comparisons [146]. Complete [110]. Complex [271, 28]. Complex-Valued [271]. Compositing [128]. compressed [91]. Computation [144, 218, 187, 52, 103, 95, 99, 8]. Computational [286]. Computer [16, 24]. Computer-Assisted [24]. Computing [269, 167, 108, 188, 123, 49, 35]. Cone [144, 86]. Cone-Spheres [144]. Cones [144]. Conformal [245]. Conservative [166]. Consistent [90]. Consolidated [312]. Constancy [236]. Constant [100]. Construction [301, 243]. Content [150]. Continuous [216, 257]. Contour [91]. Control [76]. Controllable [96]. Controlled [153]. Controls [16]. Conventional [51]. Conversion [226, 59]. Convex [204, 117, 77, 229, 50, 5, 90]. Convolution [89]. Convolutions [259]. Coordinate [293]. Coordinates [167, 297, 68, 142, 98, 138, 3]. correct [13]. Counterexample [297]. Creases [149]. Creating [14]. Crooked [126]. Crossing [153]. Cube [216, 127]. Cubes [265, 3]. Cubic [302, 294, 288, 111]. Cueing [121]. Culling [64, 65, 32, 266]. Cursor [26]. Cursors [109]. Curve [228, 108, 134, 95]. Curved [311]. Curves [191, 181, 289, 288, 111, 282, 83].

Curvilinear [207]. Cyclic [293]. Cylinders [144]. Cylindrical [116, 290, 3].

D [4, 82, 237, 168, 79, 78, 114, 291, 181, 154, 312, 295, 109, 183, 99]. Damped [171]. Dance [313]. Data [309, 163]. DE-Tree [192]. Debugging [105]. Decide [77]. Decorative [225]. Deferred [56]. Defined [188]. Deformable [304, 299, 28]. Deformation [40, 63, 68, 201, 213]. Deformations [174]. Deforming [62, 218, 23]. Degree [136, 149]. Delaunay [299]. Demosaic [234]. Dependent [257]. Depth [221, 194, 43, 58, 266, 187, 169, 276, 121]. Depth-Based [169]. Depth-Buffer-Based [58]. Depth-Buffering [43]. Depth-Cueing [121]. Derivation [51]. Descent [293]. Detail [223, 257]. Detection [60, 47, 195, 248, 143, 19, 111, 28, 50]. Determining [54]. Device [76]. Diagrams [196, 4]. Dielectrics [102]. Differentiation [152]. Diffusion [154, 193]. Diffusion-based [154]. Dimensional [25, 44, 68, 155, 76, 162, 268, 52, 77, 111, 72, 51, 92]. Dimensionality [307]. Dimensions [76]. Direct [292, 185]. Directed [42]. Directional [68]. Disc [245]. Discrepancy [287]. Discrete [289]. Disk [260, 186, 270, 27]. Display [309, 306]. Displaying [208]. Distance [144, 196, 305, 310, 206, 181, 257, 99]. Distance-Dependent [257]. Distances [124]. Distortion [27]. Distributed [100]. Distribution [260]. distributions [89]. Divergence [177]. Divergence-Free [177]. Divisions [118]. Dome [255]. Drawing [298]. Dynamic [250, 33, 265, 242, 145, 128]. Dynamics [304].

easy [86]. Edge [164]. Edges [42, 229]. Edging [45]. Editorial [279, 296, 314, 308, 273]. Editors [303].

Efficiency [36, 199]. Efficient

[302, 96, 164, 180, 294, 107, 228, 30, 186, 55, 175, 222, 127, 286, 234, 275, 238, 170, 243, 282, 158, 131, 28, 6]. Efficiently [57]. Elastic [38]. Ellipse [140, 92]. Ellipsoid [222]. Ellipsoid-OBB [222]. Elliptical [264]. Empty [227]. Empty-Space [227]. Enclosing [305]. Encoded [84]. Encoding [25, 33]. Energy [300]. Enhanced [226, 94]. Enumeration [164]. Environment [237, 180, 246, 208, 112]. EOV [279, 296, 314]. Equal [233]. Equal-Area [233]. ERIT [30]. Erratum [302]. Estimation [101]. Evaluation [211, 161, 271]. Exact [211, 183]. Expenditure [300]. Exponential [41, 271]. Exposures [128]. Extended [74]. extraction [91]. Extrusion [154].

Facet [49]. Facets [35]. Factor [93].

Factorization [119]. Factors [178]. False [31]. Fast [82, 144, 44, 168, 233, 269, 217, 196, 299, 104, 125, 9, 29, 58, 142, 209, 8, 17, 21, 93, 218, 138, 187, 163, 52, 143, 124, 245, 103, 141, 77, 95, 111, 99, 120, 147, 128, 86, 122, 50, 7]. Faster [53, 305, 198, 212]. Features [311]. FFD [213]. FFT [87]. fi [7]. Field [201, 266]. Fields [177, 84]. Filling [282]. Filtering [305, 234, 157]. Filters [156, 244]. First [94]. Fitting [228]. Flattening [240]. Flexible [223]. Floater [288]. fluid [87]. Fog [206]. Fogging [121]. Foliage [179]. Form [40, 63, 178, 213, 93, 111]. Formulas [206]. Forward [263, 72]. Four [54]. Fourier [284]. Framework [223, 161]. Frédéric [302]. Free [202, 40, 63, 213, 177, 253]. Free-Form [40, 63, 213]. Free-Viewpoint [253]. Frequency [258, 285]. Frustum [64, 123]. Full [33]. Full-Gamut [33]. Function [96]. Functions [72].

Gamut [33]. gaseous [88]. Gathering [22].

General [203]. Generalized [297, 98, 134]. Generating [191, 75, 83]. Generation [160, 24, 186, 157, 120]. Geometric [105, 231, 111, 23]. Geometry [159]. GJK [50]. Global [53, 165, 272, 12]. GLSL [286]. Golden [287]. Goniophotometric [146]. GPU [302, 241, 194, 259, 294, 274, 236, 196, 304, 305, 201, 261, 203, 288, 235, 238]. GPU-Based [241, 194, 274, 304, 201, 238]. GPUs [234, 165]. Graph [302]. Graphics [70, 202, 178, 93, 303, 174, 131, 122, 51]. Graphs [67]. Gray [226]. Grid [299, 298]. Grids [207]. grouping [79]. Guarantees [127].

Halfspace [206]. Halftoning [225, 115]. Halton [20]. Hammersley [20]. Hand [128]. Hand-Held [128]. Handling [176, 195]. Hardware [70, 202, 160, 159, 114, 117, 116, 178, 88, 93, 289, 174, 131, 122, 18]. Hardware-Accelerated [160, 88]. Hardware-Assisted [159, 116]. Harmonic [259, 185]. Hashing [248]. having [5]. Head [208]. Head-Tracked [208]. Heightmaps [257]. Held [128]. Hemi [233]. Hemisphere [100]. Heuristic [31]. Hexagonal [298]. Hexahedron [207]. Hi [7]. hi-fi [7]. hi-lights [7]. Hi-speed [7]. Hierarchical [251, 178, 248]. Hierarchies [267]. High [258, 24, 196, 33, 181, 242, 234, 18, 128]. High-Dynamic [33]. High-Frequency [258]. High-Order [196]. High-Performance [181]. High-Quality [234]. High-Reliefs [24]. Hilbert [134]. hue [1]. hue-based [1]. Hull [77]. HWB [1].

IBar [172]. Icons [154]. Illumination [53, 165, 12, 7]. Image [302, 205, 294, 123, 34, 135, 141, 128, 220, 122, 51, 39]. Image-Based [135, 51]. Image-Space [220]. Images [33, 242, 154, 102, 97, 91]. Implementation [304, 127, 203, 50, 12]. Implementing [240, 172]. implicit [6]. Implicitization [288]. Impostors [160].

Improved [182, 310, 298, 163, 312]. Improving [4, 129, 146, 199]. Inclusion [77]. Incremental [260, 105, 62, 210]. Indexing [232]. Inertia [188]. Inexpensive [39, 255]. Inflatable [154]. Information [25, 80]. Inpainting [141]. Input [76]. Integer [256]. Integral [205, 271]. Integration [183]. Intelligent [67]. Interaction [109]. Interactive [184, 250, 304, 179, 154, 252, 227, 106]. Interference [195]. Interpolation [302, 294, 43, 238, 46]. Interpolations [162]. Intersection [168, 30, 47, 175, 222, 161, 17, 21, 138, 158]. Intersections [148]. Intrusive [105]. intuitive [1]. Inverse [171, 293, 72]. Irradiance [280]. Irregular [98]. iSlerp [210]. Isolation [192]. Isophotes [10]. Issue [303]. Issues [36].

J [302]. Joint [239, 86]. Journal [303].

JPEG [91].

Kinematics [171, 293, 254].

Large [309]. Laser [130]. Layout [282]. Least [171]. Length [95]. Level [223, 257]. Library [145]. Light [246, 255, 84]. Lighting [16, 190]. lights [7]. Lightweight [197]. limits [86]. Line [144, 168, 281, 295, 99, 229, 89, 18]. Line-Swept [144]. Linear [79, 270, 244]. Linear-Time [270]. Linearly [218]. Lines [54, 13]. List [137]. List-Priority [137]. Local [187]. LogLuv [33]. Low [9, 285, 287, 27]. Low-Discrepancy [287]. Low-Frequency [285].

Management [150]. Managing [272, 229]. Manipulation [292, 240]. Map [53, 246, 41, 27, 120]. Mapping [263, 132, 221, 233, 216, 285, 214, 261, 212, 245, 72, 113]. Mappings [297]. Maps [202, 180, 173, 189, 112]. Marching [3, 265, 127, 141]. Mass [237, 8]. Matrix [57]. Maximize [123]. Maya [153]. Maya^{T M} [150]. Medium [9]. Medium-Quality [9]. Membranes [38]. Memory [9]. Mesh [188, 185, 145, 282, 23]. Meshes [42, 299, 130, 195, 218, 170, 38, 73, 120, 80]. Message [303]. Method [274, 181, 300, 187, 212, 141, 48, 45, 79]. Methods [107, 19]. Metric [151]. Minimum [305, 181, 17, 300]. Mipmapping [173]. Mirror [306]. Mix [254]. Model [133, 69, 182, 1, 7]. Models [311, 154, 193, 23, 28, 14]. Modern [131]. Modified [166]. Moment [188]. Monitor [306]. Motion [266, 199]. MPEG [84]. MPEG-Encoded [84]. Multi [284, 252, 48]. Multi-Band [284]. Multi-Stage [48]. Multi-Touch [252]. multiple [14]. Multiresolution [265, 201].

Near [211]. Nearest [248]. Neighbor [241, 248]. Neighbor-Search [241]. Neighborhood [232]. Nested [102]. Newell [103]. Night [97]. Noise [286, 220]. Non [105]. Non-Intrusive [105]. Nonmanifold [230, 231]. Nonoriented [231]. Normal [285, 201, 103, 173, 120]. Normalization [182, 291]. Normals [49, 218, 35]. NURBS [62, 181, 66].

OBB [222, 143]. Object [123, 192]. Objects [44, 40, 63, 312, 19, 50]. Observations [139]. Obtain [39]. Occlusion [209]. Ocean [284]. Octrees [107]. One [34, 57, 244, 72]. One-Dimensional [72]. One-Pass [34]. Opaque [230]. Operators [145]. Optimal [156]. Optimized [132, 64, 46]. optimizing [79]. Order [196]. Orientation [125]. Orthonormal [291, 61]. Overlap [217, 104, 125, 198, 142, 124, 82].

Packing [135]. Panoramas [116].

Parallelization [305]. Parameter [101]. Parameterization [200, 41]. Parameterizations [290]. Parametric [191, 95, 92]. Particle [241, 250, 47]. Pass [129, 34]. Patch [148]. Path [300, 247]. Paths [246]. Patterns [263, 186, 270]. PB [213]. PB-FFD [213]. Peeling [221, 194]. Axis-Aligned [217]. Cost [146]. Penetration [187]. Penumbra [189]. Per-Pixel [118]. Perceptual [151]. Performance [181]. Performing [9]. Perpendicular [243]. Perspective [214, 283]. Phase [248]. phenomena [88]. Phong [69, 7]. Photogrammetry [152]. Photographcs [128]. Photographic [101]. Photon [53]. Photorealistic [153]. Physical [254]. Physics [307]. Pixel [263, 118, 157]. planar [81, 80]. Plane [246, 267, 214, 26]. plateaus [81]. Plücker [142, 138]. Point [160, 144, 167, 117, 213, 268, 170, 77]. Point-Based [160, 213]. Points [75, 100, 20]. Poisson [260, 186, 270]. Poisson-Disk [260, 186, 270]. Polar [249, 68]. Polygon [164, 103, 45, 229]. Polygonal [40, 63, 110, 35, 23]. Polygonization [62, 6]. Polygons [137, 155, 162, 98, 5]. polyhedra [90]. polyhedral [8]. Polylines [229]. Polynomial [156, 89]. Polytopes [117]. Position [304]. Possible [240]. Post [97]. Post-Processing [97]. Practical [292, 113, 137, 41, 66]. Precision [283]. Precomputed [285, 190, 209]. Predicates [125]. Prediction [307]. Prefiltering [155]. Price [244]. Priority [137]. Procedural [133, 310]. Processing [97]. Product [262]. Production [151]. Programmable [178]. Progressive [195]. Projected [167, 52]. Projection [207]. properties [8]. Proxies [135]. Proximity [224, 209]. Pseudo [109]. Pseudo-Shadowed [109]. Pushing [153].

Quadrilateral [175, 275, 183].

Quadrilaterals [204]. Quadtrees [107]. Quality [9, 234, 18]. Quantization [263]. Quaternions [301]. Queries [70, 100, 192]. Query [232, 117].

Radiosity [22, 178, 93, 19, 276]. Rainfall [235]. Random [75]. Range [33, 242, 128]. Rapid [195]. Rapidly [10]. Rasterization [166, 117]. Ratio [146, 287]. Rational [191, 83]. Ray [184, 194, 197, 217, 267, 29, 256, 310, 175, 161, 142, 66, 17, 94, 138, 148, 102, 15, 36, 46, 158, 192, 106]. Ray-Axis [142]. Ray-Box [158]. Ray-Quadrilateral [175]. Ray-Tetrahedron [138]. Ray-Triangle [17]. Ray/Axis [217]. Ray/ Axis-Aligned [217]. reach [86]. reach-cone [86]. Reaction [193]. Reaction-Diffusion [193]. Readable [208]. Real [258, 78, 251, 115, 135, 165, 119, 121, 174, 254, 88]. Real-Time [258, 251, 115, 135, 165, 119, 121, 174, 254, 78, 88]. Realistic [147]. Recipe [162]. Reconciling [264]. Reconstruction [201]. Recursive [71]. Reducing [220]. Reduction [307]. Reflectance [182]. Reflectances [59]. reflections [112]. Reflectivity [274]. Region [130]. Region-Based [130]. Registration [128]. Regression [307]. Regular [73]. Relabeling [299]. Reliability [146]. Reliability/Cost [146]. Reliable [30]. Relief [212]. Reliefs [24]. Removal [227]. Removing [249]. Renderer [253]. Rendering [249, 160, 159, 258, 239, 22, 250, 9, 116, 179, 135, 289, 288, 119, 257, 283, 227, 147, 51, 78, 88]. RenderMan [12, 153]. Representation [42, 231, 80]. Reproduction [101, 272]. Requests [232]. Resampling [129, 34, 72]. RGB [59]. RGB-to-Spectrum [59]. Rigid [240]. Rigs [280]. Robust [221, 125, 29, 128, 158, 50, 4]. rooks [48]. Rotate [57]. Rotations [41].

Sampled [180]. Samples [157]. Sampling [204, 246, 186, 190, 287, 134, 199, 48, 140, 20]. Sant [302]. Scalable [42, 203]. Scanned [130]. Scene [251, 110]. Scenes [97]. Scheme [248]. Scientific [39]. Scissors [67]. Screen [250]. Screen-Space [250]. Seams [153, 290]. Search [241, 248]. Secant [212]. Segment [168]. Segmentation [122]. Selection [295]. Selective [169, 15]. Selectively [171]. Self [114]. Self-Shadowing [114]. Semi [73]. Semi-Regular [73]. Separating [214]. Separating-Plane [214]. Sequences [287]. Set [166, 77]. Set-Up [166]. Sets [268]. Shaded [45]. Shader [312]. Shaders [275]. Shading [277, 278]. Shadow [136, 132, 202, 221, 223, 219, 55, 230, 214, 19, 189, 113]. Shadow-Volume [230]. Shadowed [109]. Shadowing [114]. Shadows [180, 258, 223, 209, 81]. Shaft [65]. Shape [240]. Shared [267]. Shared-Plane [267]. Sharp [311, 19]. Shortest [300]. Shot [150]. Shows [313]. Shutter [199]. Signed [124]. SIGRAD [303]. Silhouette [136, 139]. SIMD [233]. Simple [112, 71, 107, 177, 102, 272, 6, 227, 282, 31, 87]. Simplification [130]. Simplified [120]. Simulation [307, 177, 38, 254]. Simulations [241]. Single [39, 306]. Single-image [39]. Single-Monitor-Mirror [306]. Sized [220]. Sizes [139]. Skeleton [23]. skeletons [89]. SLERP [269, 210]. Slice [159]. Slicing [60]. Slicing-Based [60]. Slimmed [67]. Slopes [217]. Small [301]. Smooth [162, 189]. Smoothing [122]. Soft [221, 81]. Solid [188]. Solutions [22]. solver [87]. Space [153, 250, 227, 282, 220]. Space-Filling [282]. Spatial [97]. Special [303]. Spectral [268]. Spectrum [59]. specular [7]. Speed [70, 7]. Sphere [232, 233, 198, 301]. Sphere-Box [198]. Spheres [144, 14]. Spherical [259, 200, 185, 3]. Splat [311]. Splat-based [311]. Splatting [220, 131]. Spline [211, 157, 37, 302, 294, 40, 63]. Splines [238]. Spring [38]. Square [27, 245]. Squares [171]. Stable [245]. Stage [255, 48]. Stair [85]. standard [12]. Star [80]. Star-vertices [80]. Stellar [145]. Stencil [132]. Stereoscopic [208, 306]. Stereovision [39]. Stethoscope [105]. Stippling [260]. Stochastic [179, 276]. Storage [17]. Stratified [204, 157, 134, 140]. Stretch [262]. Stripping [163]. Structure [163]. Subdivision [249, 211, 149, 215, 73, 90, 83]. Subtractive [223]. Surface [25, 71, 250, 240, 135]. Surfaces [249, 184, 211, 62, 66, 261, 252, 73, 45, 37, 89, 6]. Swept [144]. Synthesis [284, 193]. Synthetic [44]. System [150]. Systems [47].

Technique [213, 19, 141, 140]. Techniques [184]. Templates [152]. Temporal [199]. Temporally [106]. Term [74, 7]. Tessellation [289]. Tessellator [71]. Test [168, 125, 47, 175, 222, 21, 124]. TESTIMAGES [309]. Testing [309, 198, 151, 82]. Tests [217, 30, 142]. tetrahedra [90]. Tetrahedral [299]. Tetrahedron [104, 138, 75]. Tetrahedron-Tetrahedron [104]. Text [208]. Texture [133, 153, 25, 114, 135, 93, 238, 193, 262]. Texture-Based [93]. Textured [219]. Texturing [37]. Their [244]. Thread [133]. Three [25, 44, 68, 76, 52, 51, 92]. Three-Dimensional [25, 44, 68, 52, 51, 92]. Thresholding [205]. Tight [108]. Tightening [283]. Tiled [166, 194, 277]. Time [258, 251, 270, 115, 135, 165, 100, 119, 121, 174, 254, 78, 88]. Tone [101, 272]. Tool [65, 313, 295]. Tools [302, 303]. Topological [127]. Toroidal [290]. Touch [252]. Traced [102]. Tracing [184, 197, 267, 29, 256, 66, 94, 247, 36, 106, 15]. Tracked [208]. Transform [185]. transformation [79]. Transforms [259, 196]. Transition [265]. Transitions [189]. Transparent [230]. Transposes [244]. Traversal [107, 29]. Tree [29, 192]. Trees [258, 224, 195, 143, 28]. Tri [26]. Tri-Plane [26]. Triangle [166, 42, 168, 43, 125, 188, 161, 275, 17, 21, 163, 170, 124, 262, 82]. Triangle-Based [43]. triangle-box [82]. Triangle-Triangle [125, 21, 124]. Triangulated [185, 38]. Triangulating [5]. Triangulation [71]. Tricks [244]. Trimmed [66]. Tubes [289]. Twelve [13]. Two [129, 155, 76, 181, 162, 34, 244, 268, 77, 111]. Two-Dimensional [155, 76, 162, 268, 77, 111]. Two-Pass [129, 34].

Unbounded [237]. Unconstrained [200]. Unified [206]. Uniform [299, 238]. Uniformly [100]. Union [117, 14]. Unit [291, 61, 301]. Unparametrized [261]. Unstructured [239]. Update [143]. Upsampling [239]. Using [166, 70, 153, 132, 39, 194, 299, 41, 68, 55, 195, 135, 93, 289, 138, 312, 124, 134, 276, 73, 67, 131, 122, 254, 202, 260, 205, 233, 307, 217, 114, 125, 81, 117, 178, 201, 212, 238, 174, 28]. UV [153].

Valued [271]. Variable [220]. Variable-Sized [220]. Vector [25, 291, 61, 57]. Vectors [243]. Versatile [225]. Version [34]. Vertex [49, 218, 290, 35, 79]. Vertices [136, 5, 80]. Video [253]. View [64, 43, 123]. Viewpoint [253]. views [14]. Virtual [208]. Visibility [70, 285, 15]. Visualization [39, 105]. Visualizing [10, 242]. Volume [136, 159, 239, 267, 230, 227, 46]. Volumes [197, 219]. Volumetric [60, 190]. Voronoi [196, 4]. Voxel [176]. Voxel-Based [176]. Voxelization [44, 251, 110, 58, 203].

Wallpaper [126]. Ward [182]. Waves [284, 31]. Web [242]. weight [89]. Weights [49]. Wide [192]. Widget [172]. Within [275]. Without [291, 290, 43, 155, 243]. Wrap [278].

Yves [302].

References

Smith:1996:HMI

[1] Alvy Ray Smith and Eric Ray Lyons. HWB: A more intuitive hue-based color model. Journal of Graphics Tools: JGT, 1(1):3–17, 1996. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/SmithLyons96/.

Bigos:1996:ABC

[2] Andrew Bigos. Avoiding buffer clears. Journal of Graphics Tools: JGT, 1(1): 19–20, 1996. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/Bigos96/.

Goldsmith:1996:MCC

[3] Jeff Goldsmith and Allan S. Jacobson. Marching cubes in cylindrical and spherical coordinates. Journal of Graphics Tools: JGT, 1(1):21–32, 1996. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/GoldsmithJacobson96/.

Hubbard:1996:IAR

[4] Philip M. Hubbard. Improving accuracy in a robust algorithm for 3D Voronoi diagrams. Journal of Graphics Tools: JGT, 1(1):33–45, 1996. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/Hubbard96/.

Cignoni:1996:TCP

[5] P. Cignoni, C. Montani, and R. Scopigno. Triangulating convex polygons having T -vertices. Journal of Graphics Tools: JGT, 1(2):1–4, 1996. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/CignoniMontaniScopigno96/.

Velho:1996:SEP

[6] Luiz Velho. Simple and efficient polygonization of implicit surfaces. Journal of Graphics Tools: JGT, 1(2):5–24, 1996. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Velho96/.

vanOverveld:1996:HSH

[7] C. W. A. M. van Overveld and Brian Wyvill. Hi-speed, hi-fi hi-lights: a fast algorithm for the specular term in the Phong illumination model. Journal of Graphics Tools: JGT, 1(2):25–30, 1996. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/vanOverVeldWyvill96/.

Mirtich:1996:FAC

[8] Brian Mirtich. Fast and accurate computation of polyhedral mass properties. Journal of Graphics Tools: JGT, 1(2):31–50, 1996. CODEN JGTOFD. ISSN 1086-7651. URL http://

www.acm.org/jgt/papers/Mirtich96/
;   http://www.cs.berkeley.edu/~
mirtich/massProps.html.

Haines:1996:FLM

[9] Eric Haines and Steven Worley. Fast, low memory Z-buffering when performing medium-quality rendering. Journal of Graphics Tools: JGT, 1(3):1–5, 1996. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/HainesWorley96/.

Hutchinson:1996:RVI

[10] Dave Hutchinson and Terry Hewitt. Rapidly visualizing isophotes. Journal of Graphics Tools: JGT, 1(3):7–12, 1996. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/HutchinsonHewitt96/.

vanOverveld:1996:BBB

[11] C. W. A. M. van Overveld and Brian Wyvill. Banishing bad buckling. Journal of Graphics Tools: JGT, 1(3):13– 28, 1996. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/vanOverVeldWyvill96b/;
 http://www.win.tue.nl/win/cs/tt/
kees/PAPERS_PS/b_buckl.ps.gz.

Gritz:1996:BGI

[12] Larry Gritz and James K. Hahn. BMRT: A global illumination implementation of the RenderMan standard. Journal of Graphics Tools: JGT, 1(3):29–47, 1996. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/GritzHahn96/.

Nelson:1996:TCC

[13] Scott R. Nelson. Twelve characteristics of correct antialiased lines. Journal of Graphics Tools: JGT, 1(4):1–20, 1996. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Nelson96/.

Ranjan:1996:CUS

[14] Vishwa Ranjan and Alain Fournier. Creating union of spheres models from multiple views. Journal of Graphics Tools: JGT, 1(4):21–39, 1996. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/RanjanFournier96/.

Sherstyuk:1996:RTS

[15] Andrei Sherstyuk. Ray tracing with selective visibility. Journal of Graphics Tools: JGT, 1(4):41–46, 1996. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/Sherstyuk96/.

Barzel:1997:LCC

[16] Ronen Barzel. Lighting controls for computer cinematography. Journal of Graphics Tools: JGT, 2(1):1–20, 1997. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/Barzel97/.

Moller:1997:FMS

[17] Tomas Möller and Ben Trumbore. Fast, minimum storage ray-triangle intersection. Journal of Graphics Tools: JGT, 2(1):21–28, 1997. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/MollerTrumbore97/.

Nelson:1997:HQH

[18] Scott R. Nelson. High quality hardware line antialiasing. Journal of Graphics Tools: JGT, 2(1):29–46, 1997. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/Nelson97/.

Telea:1997:COB

[19] A. C. Telea and C. W. A. M. van Overveld. The close objects buffer: A sharp shadow detection technique for radiosity methods. Journal of Graphics Tools: JGT, 2(2):1–8, 1997. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/TeleaVanOverveld97/.

Wong:1997:SHH

[20] Tien-Tsin Wong, Wai-Shing Luk, and Pheng-Ann Heng. Sampling with Hammersley and Halton points. Journal of Graphics Tools: JGT, 2(2):9–24, 1997. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/WongLukHeng97/.

Moller:1997:FTT

[21] Tomas Möller. A fast triangletriangle intersection test. Journal of Graphics Tools: JGT, 2(2):25–30, 1997. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Moller97/.

Chung:1997:RRS

[22] A. J. Chung and A. J. Field. Rendering radiosity solutions by adaptive gathering. Journal of Graphics Tools: JGT, 2(2):31–44, 1997. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/ChungField97/.

vanOverveld:1997:DGM

[23] C. W. A. M. van Overveld and M. G. J. R. Stalpers. Deforming geometric models based on a polygonal skeleton mesh. Journal of Graphics Tools: JGT, 2(3):1–14, 1997. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/StalpersVanOverveld97/.

Cignoni:1997:CAG

[24] P. Cignoni, C. Montani, and R. Scopigno. Computer-assisted generation of basand high-reliefs. Journal of Graphics Tools: JGT, 2(3):15–28, 1997. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/CignoniMontaniScopigno97/.

Bailey:1997:ETD

[25] Michael Bailey and Dru Clark. Encoding three-dimensional surface information in a texture vector. Journal of Graphics Tools: JGT, 2(3):29–35, 1997. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/BaileyClark97/.

Xiang:1997:TPC

[26] Zhigang Xiang. A tri-plane cursor. Journal of Graphics Tools: JGT, 2(3):37–43, 1997. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/Xiang97/.

Shirley:1997:LDM

[27] Peter Shirley and Kenneth Chiu. A low distortion map between disk and square. Journal of Graphics Tools: JGT, 2(3): 45–52, 1997. CODEN JGTOFD. ISSN

                                        1086-7651. URL http://www.acm.org/
                                        jgt/papers/ShirleyChiu97/.

vandenBergen:1997:ECD

[28] Gino van den Bergen. Efficient collision detection of complex deformable models using AABB trees. Journal of Graphics Tools: JGT, 2(4):1–14, 1997. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/vanDenBergen97/.

Havran:1997:FRB

[29] Vlastimil Havran, Tomas Kopal, Jiri Bittner, and Jiri Zara. Fast robust BSP tree traversal algorithm for ray tracing. Journal of Graphics Tools: JGT, 2(4):15–24, 1997. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/HavranKopalBittnerZara97/.

Held:1997:ECE

[30] Martin Held. ERIT: A collection of efficient and reliable intersection tests. Journal of Graphics Tools: JGT, 2(4): 25–44, 1997. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/Held97/.

vanOverveld:1997:CWS

[31] C. W. A. M. van Overveld. Color waves: A simple heuristic for choosing false colors. Journal of Graphics Tools: JGT, 2(4):45–50, 1997. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/vanOverveld97/

.

Johannsen:1998:CBC

[32] Andreas Johannsen and Michael B. Carter. Clustered backface culling. Journal of Graphics Tools: JGT, 3(1): 1–14, 1998. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/JohannsenCarter98/.

Larson:1998:LEF

[33] Gregory Ward Larson. LogLuv encoding for full-gamut, high-dynamic range images. Journal of Graphics Tools: JGT, 3(1):15–31, 1998. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/Larson98/.

Max:1998:OPV

[34] Nelson Max. A one-pass version of two-pass image resampling. Journal of Graphics Tools: JGT, 3(1):33–41, 1998. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Max98/.

Thurmer:1998:CVN

[35] Grit Thürmer and Charles A. Wüthrich. Computing vertex normals from polygonal facets. Journal of Graphics Tools: JGT, 3(1):43–46, 1998. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/ThurmerWuthrich98/.

Smits:1998:EIR

[36] Brian Smits. Efficiency issues for ray tracing. Journal of Graphics Tools: JGT, 3(2):1–14, 1998. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.acm.org/jgt/papers/Smits98/

.

Woo:1998:CTS

[37] Andrew Woo. Chordlength texturing of spline surfaces. Journal of Graphics Tools: JGT, 3(2):15–19, 1998. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.acm.org/jgt/papers/Woo98/.

VanGelder:1998:ASE

[38] Allen Van Gelder. Approximate simulation of elastic membranes by triangulated spring meshes. Journal of Graphics Tools: JGT, 3(2):21–41, 1998. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/VanGelder98/.

Bailey:1998:UCO

[39] Michael Bailey and Dru Clark. Using ChromaDepth to obtain inexpensive single-image stereovision for scientific visualization. Journal of Graphics Tools: JGT, 3(3):1–9, 1998. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/BaileyClark98/.

Feng:1998:ABS

[40] Jieqing Feng, Pheng-Ann Heng, and Tien-Tsin Wong. Accurate B-spline freeform deformation of polygonal objects. Journal of Graphics Tools: JGT, 3(3): 11–27, 1998. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/FengHengWong98/.

Grassia:1998:PPR

[41] F. Sebastian Grassia. Practical parameterization of rotations using the exponential map. Journal of Graphics Tools: JGT, 3(3):29–48, 1998. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Grassia98/.

Campagna:1998:DES

[42] Swen Campagna, Leif Kobbelt, and Hans-Peter Seidel. Directed edges — A scalable representation for triangle meshes. Journal of Graphics Tools: JGT, 3(4):1–12, 1998. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/CampagnaKobbeltSeidel98/.

Fu:1998:TBV

[43] Chi-Wing Fu, Tien-Tsin Wong, and Pheng-Ann Heng. Triangle-based view interpolation without depth-buffering. Journal of Graphics Tools: JGT, 3(4): 13–31, 1998. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/FuWongHeng98/.

Chen:1998:FVT

[44] Hongsheng Chen and Shiaofen Fang. Fast voxelization of three-dimensional synthetic objects. Journal of Graphics Tools: JGT, 3(4):33–45, 1998.

CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/ChenFang98/.

Wang:1999:NMP

[45] Wencheng Wang, Yanyun Chen, and Enhua Wu. A new method for polygon edging on shaded surfaces. Journal of Graphics Tools: JGT, 4(1):1–10, 1999. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/WangChenWu99/.

Wan:1999:OIV

[46] Ming Wan, Arie Kaufman, and Steve Bryson. Optimized interpolation for volume ray casting. Journal of Graphics Tools: JGT, 4(1):11–24, 1999. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/WanKaufmanBryson99/.

Karabassi:1999:ITC

[47] Evaggelia-Aggeliki Karabassi, Georgios Papaioannou, and Theoharis Theoharis. Intersection test for collision detection in particle systems. Journal of Graphics Tools: JGT, 4(1):25–37, 1999. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/KarabassiEtAl99/.

Wang:1999:MSR

[48] Changyaw Wang and Kelvin Sung. Multi-stage N-rooks sampling method. Journal of Graphics Tools: JGT, 4(1): 39–47, 1999. CODEN JGTOFD. ISSN

                                        1086-7651. URL http://www.acm.org/
                                        jgt/papers/WangSung99/.

Max:1999:WCV

[49] Nelson Max. Weights for computing vertex normals from facet normals. Journal of Graphics Tools: JGT, 4(2):1–6, 1999. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Max99/.

vandenBergen:1999:FRG

[50] Gino van den Bergen. A fast and robust GJK implementation for collision detection of convex objects. Journal of Graphics Tools: JGT, 4(2):7–25, 1999. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/vanDenBergen99/.

Zhang:1999:DIB

[51] Hansong Zhang. A derivation of image-based rendering for conventional three-dimensional graphics. Journal of Graphics Tools: JGT, 4(2):27–36, 1999. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Zhang99/.

Schmalstieg:1999:FPA

[52] Dieter Schmalstieg and Robert F. Tobler. Fast projected area computation for three-dimensional bounding boxes. Journal of Graphics Tools: JGT, 4(2): 37–43, 1999. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/SchmalstiegTobler99/.

Christensen:1999:FPM

[53] Per H. Christensen. Faster photon map global illumination. Journal of Graphics Tools: JGT, 4(3):1–10, 1999. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Christensen99/.

Teller:1999:DLT

[54] Seth Teller and Michael Hohmeyer. Determining the lines through four lines.

Journal of Graphics Tools: JGT, 4(3): 11–22, 1999. CODEN JGTOFD. ISSN

                                        1086-7651. URL http://www.acm.org/
                                        jgt/papers/TellerHohmeyer99/.

Keating:1999:ESA

[55] Brett Keating. Efficient shadow antialiasing using an A-buffer. Journal of Graphics Tools: JGT, 4(3):23–33, 1999. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/Keating99/.

McCormick:1999:DAB

[56] Patrick S. McCormick, Charles Hansen, and Edward Angel. The deferred accumulation buffer. Journal of Graphics Tools: JGT, 4(3):35–46, 1999. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/McCormickHansenAngel99/.

Moller:1999:EBM

[57] Tomas Möller and John F. Hughes. Efficiently building a matrix to rotate one vector to another. Journal of Graphics Tools: JGT, 4(4):1–4, 1999. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/MollerHughes99/.

Karabassi:1999:FDB

[58] Evaggelia-Aggeliki Karabassi, Georgios Papaioannou, and Theoharis Theoharis. A fast depth-buffer-based voxelization algorithm. Journal of Graphics Tools: JGT, 4(4):5–10, 1999. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/KarabassiEtAl99b/.

Smits:1999:RSC

[59] Brian Smits. An RGB-to-spectrum conversion for reflectances. Journal of Graphics Tools: JGT, 4(4):11–22, 1999. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Smits99/.

Boyles:1999:SBV

[60] Michael Boyles and Shiaofen Fang. Slicing-based volumetric collision detection. Journal of Graphics Tools: JGT, 4(4):23–32, 1999. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/BoylesFang99/.

Hughes:1999:BOB

[61] John F. Hughes and Tomas Möller. Building an orthonormal basis from a unit vector. Journal of Graphics Tools: JGT, 4(4):33–35, 1999. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/HughesMoller99/.

Li:1999:IPD

[62] Frederick W. B. Li and Rynson W. H. Lau. Incremental polygonization of deforming NURBS surfaces. Journal of Graphics Tools: JGT, 4(4):37–50, 1999. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/LiLau99/.

Feng:2000:AAB

[63] Jieqing Feng and Qunsheng Peng. Accelerating accurate B-spline free-form deformation of polygonal objects. Journal of Graphics Tools: JGT, 5(1):1–8, 2000. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/FengPeng00/.

Assarsson:2000:OVF

[64] Ulf Assarsson and Tomas Möller. Optimized view frustum culling algorithms for bounding boxes. Journal of Graphics Tools: JGT, 5(1):9–22, 2000. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/AssarssonMoller00/.

Haines:2000:SCT

[65] Eric Haines. A shaft culling tool. Journal of Graphics Tools: JGT, 5(1):23–26, 2000. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/Haines00/.

Martin:2000:PRT

[66] William Martin, Elaine Cohen, Russell Fish, and Peter Shirley. Practical ray tracing of trimmed NURBS surfaces. Journal of Graphics Tools: JGT, 5(1): 27–52, 2000. CODEN JGTOFD. ISSN

                                        1086-7651. URL http://www.acm.org/
                                        jgt/papers/MartinCohenFishShirley00/

.

Wong:2000:AIS

[67] Kevin Chun-Ho Wong, Pheng-Ann Heng, and Tien-Tsin Wong. Accelerating “intelligent scissors” using slimmed graphs. Journal of Graphics Tools: JGT, 5(2):1–13, 2000. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/WongHengWong00/.

Jin:2000:TDD

[68] Xiaogang Jin and Y. F. Li. Threedimensional deformation using directional polar coordinates. Journal of Graphics Tools: JGT, 5(2):15–24, 2000. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/JinLi00/.

Ashikhmin:2000:APB

[69] Michael Ashikhmin and Peter Shirley. An anisotropic Phong BRDF model. Journal of Graphics Tools: JGT, 5(2): 25–32, 2000. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/AshikhminShirley00/.

Alonso:2000:UGH

[70] Laurent Alonso and Nicolas Holzschuch. Using graphics hardware to speed up your visibility queries. Journal of Graphics Tools: JGT, 5(2):33–47, 2000. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/AlonsoHolzschuch00/.

Chung:2000:SRT

[71] A. J. Chung and A. J. Field. A simple recursive tessellator for adaptive surface triangulation. Journal of Graphics Tools: JGT, 5(3):1–9, 2000. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/AlonsoHolzschuch00/.

Wolberg:2000:ODR

[72] George Wolberg, H. M. Sueyllam, M. A. Ismail, and K. M. Ahmed. Onedimensional resampling with inverse and forward mapping functions. Journal of Graphics Tools: JGT, 5(3):11–33, 2000. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/AlonsoHolzschuch00/.

Velho:2000:USR

[73] Luiz Velho. Using semi-regular 4-8 meshes for subdivision surfaces. Journal of Graphics Tools: JGT, 5(3):35–47, 2000. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/AlonsoHolzschuch00/.

Castro:2000:EAT

[74] Francesc Castro, László Neumann, and Mateu Sbert. Extended ambient term. Journal of Graphics Tools: JGT, 5(4):1– 7, 2000. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/AlonsoHolzschuch00/.

Rocchini:2000:GRP

[75] C. Rocchini and P. Cignoni. Generating random points in a tetrahedron. Journal of Graphics Tools: JGT, 5(4):9–12, 2000. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/AlonsoHolzschuch00/.

Livingston:2000:CCT

[76] Mark A. Livingston, Arthur Gregory, and W. Bruce Culbertson. Camera control in three dimensions with a two-dimensional input device. Journal of Graphics Tools: JGT, 5(4):13–24, 2000. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/AlonsoHolzschuch00/.

Torres:2000:FAD

[77] J. C. Torres and F. A. Conde. A fast algorithm to decide the inclusion of a point in the convex hull of a two-dimensional point set. Journal of Graphics Tools: JGT, 5(4):25–32, 2000. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/AlonsoHolzschuch00/.

Elinas:2000:RTR

[78] Pantelis Elinas and Wolfgang Stuerzlinger. Real-time rendering of 3D clouds. Journal of Graphics Tools: JGT, 5(4):33–45, 2000. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/ElinasStuerzlinger00.html.

Dinerstein:2001:LGM

[79] Jonathan Dinerstein, Larre Egbert, and Nick Flann. Linear grouping— a method for optimizing 3D vertex transformation and clipping. Journal of Graphics Tools: JGT, 6(1):1–6, 2001. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/DinersteinEgbertFlann01.
html.

Kallmann:2001:SVC

[80] Marcelo Kallmann and Daniel Thalmann. Star-vertices: A compact representation for planar meshes with adjacency information. Journal of Graphics Tools: JGT, 6(1):7–18, 2001. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/KallmannThalmann01.html.

Haines:2001:SPS

[81] Eric Haines. Soft planar shadows using plateaus. Journal of Graphics Tools: JGT, 6(1):19–27, 2001. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Haines01.html.

Akenine-Moller:2001:FTB

[82] Tomas Akenine-Möller. Fast 3D triangle-box overlap testing. Journal of Graphics Tools: JGT, 6(1):29–33, 2001. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/AkenineMoller01.html.

Nasri:2001:SAG

[83] Ahmed Nasri and Gerald Farin. A subdivision algorithm for generating rational curves. Journal of Graphics Tools: JGT, 6(1):35–47, 2001. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

                                        acm.org/jgt/papers/.

vanderLinden:2001:MEL

[84] Jarno van der Linden and Richard Lobb. MPEG-encoded light fields. Journal of Graphics Tools: JGT, 6(2):1–15, 2001. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/vanDerLindenLobb01.html.

Molla:2001:SA

[85] Ramón Mollá and Roberto Vivó. The stair algorithm. Journal of Graphics Tools: JGT, 6(2):17–25, 2001. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/MollaVivo01.html.

Wilhelms:2001:FER

[86] Jane Wilhelms and Allen Van Gelder. Fast and easy reach-cone joint limits. Journal of Graphics Tools: JGT, 6(2): 27–41, 2001. CODEN JGTOFD. ISSN

                                        1086-7651. URL http://www.acm.org/
                                        jgt/papers/WilhelmsVanGelder01.
                                        html.

Stam:2001:SFS

[87] Jos Stam. A simple fluid solver based on the FFT. Journal of Graphics Tools: JGT, 6(2):43–52, 2001. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.acm.org/jgt/papers/Stam01.
                                        html.

Mech:2001:HAR

[88] Radomír Mech. Hardware-accelerated real-time rendering of gaseous phenomena. Journal of Graphics Tools: JGT, 6(3):1–16, 2001. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/Mech01/.

Jin:2001:CSL

[89] Xiaogang Jin, Chiew-Lan Tai, Jieqing Feng, and Qunsheng Peng. Convolution surfaces for line skeletons with polynomial weight distributions. Journal of Graphics Tools: JGT, 6(3):17–28, 2001. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/JinEtAl01/.

Max:2001:CSC

[90] Nelson Max. Consistent subdivision of convex polyhedra into tetrahedra. Journal of Graphics Tools: JGT, 6(3):29–36, 2001. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Max01/.

Wiseman:2001:CEC

[91] Yair Wiseman and Erick Fredj. Contour extraction of compressed JPEG images. Journal of Graphics Tools: JGT, 6(3): 37–43, 2001. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/WisemanFredj01/.

Andersen:2001:TDP

[92] Clark R. Andersen and William L. Buford, Jr. The three-dimensional parametric ellipse. Journal of Graphics Tools: JGT, 6(3):45–48, 2001. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/AndersenBuford01/.

Nielsen:2002:FTB

[93] Kasper Høy Nielsen and Niels Jørgen Christensen. Fast texture-based form factor calculations for radiosity using graphics hardware. Journal of Graphics Tools: JGT, 6(4):1–12, 2002. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/NielsenChristensen01/.

Nakamaru:2002:EBF

[94] Koji Nakamaru and Yoshio Ohno. Enhanced breadth-first ray tracing. Journal of Graphics Tools: JGT, 6(4):13–28, 2002. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/NakamaruOhno01/.

Vincent:2002:FAP

[95] Stephen Vincent and David Forsey. Fast and accurate parametric curve length computation. Journal of Graphics Tools: JGT, 6(4):29–40, 2002. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/VincentForsey01/.

Baranoski:2002:ECB

[96] Gladimir V. G. Baranoski and Jon Rokne. An efficient and controllable blob function. Journal of Graphics Tools: JGT, 6(4):41–54, 2002. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/BaranoskiRokne01/.

Thomspon:2002:SPP

[97] William B. Thomspon, Peter Shirley, and James A. Ferwerda. A spatial post-processing algorithm for images of night scenes. Journal of Graphics Tools: JGT, 7(1):1–12, 2002.

CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/ThompsonShirleyFerwerda02/

.

Meyer:2002:GBC

[98] Mark Meyer, Haeyoung Lee, Alan Barr, and Mathieu Desbrun. Generalized barycentric coordinates on irregular polygons. Journal of Graphics Tools: JGT, 7(1):13–22, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/MeyerEtAl02/.

Vranek:2002:FAC

[99] David Vranek. Fast and accurate circlecircle and circle-line 3D distance computation. Journal of Graphics Tools: JGT, 7(1):23–32, 2002. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/Vranek02/.

Slater:2002:CTQ

[100] Mel Slater. Constant time queries on uniformly distributed points on a hemisphere. Journal of Graphics Tools: JGT, 7(1):33–44, 2002. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/Slater02/.

Reinhard:2002:PEP

[101] Erik Reinhard. Parameter estimation for photographic tone reproduction. Journal of Graphics Tools: JGT, 7(1):45–52, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Reinhard02/.

Schmidt:2002:SND

[102] Charles M. Schmidt and Brian Budge.

Simple nested dielectrics in ray traced images. Journal of Graphics Tools:

JGT, 7(2):1–8, 2002. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.acm.org/jgt/papers/SchmidtBudge02/

.

Sunday:2002:FPA

[103] Daniel Sunday. Fast polygon area and Newell normal computation. Journal of Graphics Tools: JGT, 7(2):9–13 (??), 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Sunday02/.

Ganovelli:2002:FTT

[104] Fabio Ganovelli, Frederico Ponchio, and Claudio Rocchini. Fast tetrahedrontetrahedron overlap algorithm. Journal of Graphics Tools: JGT, 7(2):17–26, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/GanovelliPonchioRocchini02/

.

Cazals:2002:NID

[105] Frédéric Cazals. Non-intrusive debugging and incremental visualization with the geometric stethoscope. Journal of Graphics Tools: JGT, 7(2):27–40, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Cazals02/.

Martin:2002:TCI

[106] William Martin, Peter Shirley, Steven Parker, William Thompson, and Erik Reinhard. Temporally coherent interactive ray tracing. Journal of Graphics Tools: JGT, 7(2):41–48, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/MartinEtAl02/.

Frisken:2002:SET

[107] Sarah F. Frisken and Ronald N. Perry.

Simple and efficient traversal methods for quadtrees and octrees. Journal of Graphics Tools: JGT, 7(3):1–11, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/FriskenPerry02/.

Kallay:2002:CTB

[108] Michael Kallay. Computing tight bounds for a Bézier curve. Journal of Graphics Tools: JGT, 7(3):13–17, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Kallay02/.

Steed:2002:PSC

[109] Anthony Steed. Pseudo-shadowed cursors for 3D interaction. Journal of Graphics Tools: JGT, 7(3):19–25, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Steed02/.

Haumont:2002:CPS

[110] Denis Haumont and Nadine Warzée.

Complete polygonal scene voxelization. Journal of Graphics Tools: JGT, 7(3): 27–41, 2002. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/HaumontWarzee02/.

Vincent:2002:FDG

[111] Stephen Vincent. Fast detection of the geometric form of two-dimensional cubic Bézier curves. Journal of Graphics Tools: JGT, 7(3):43–51, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Vincent02/.

Ashikhmin:2002:SBR

[112] Michael Ashikhmin and Abhijeet Ghosh. Simple blurry reflections with environment maps. Journal of Graphics Tools: JGT, 7(4):3–8, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/AshikhminGhosh02/.

Brabec:2002:PSM

[113] Stefan Brabec, Thomas Annen, and Hans-Peter Seidel. Practical shadow mapping. Journal of Graphics Tools: JGT, 7(4):9–18, 2002. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.acm.org/jgt/papers/BrabecAnnenSeidel02/

.

Forsyth:2002:SSB

[114] Tom Forsyth. Self-shadowing bumpmap using 3D texture hardware. Journal of Graphics Tools: JGT, 7(4):19–26, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Forsyth02/.

Kautz:2002:RTH

[115] Jan Kautz and Hans-Peter Seidel. Realtime halftoning. Journal of Graphics Tools: JGT, 7(4):27–31, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/KautzSeidel02/.

Kim:2002:HAR

[116] Dongo Kim and James K. Hahn. Hardware-assisted rendering of cylindrical panoramas. Journal of Graphics Tools: JGT, 7(4):33–42, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/KimHahn02/.

Kim:2002:CPQ

[117] Young J. Kim, Kenneth Hoff, Ming C.

Lin, and Dinesh Manocha. Closest point query among the union of convex polytopes using rasterization hardware. Journal of Graphics Tools: JGT, 7(4): 43–51, 2002. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/KimEtAl02/.

Parilov:2002:PPD

[118] Sergey Parilov and Wolfgang Stuerzlinger. Per-pixel divisions. Journal of Graphics Tools: JGT, 7(4):53–59, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/ParilovStuerzlinger02/.

Steigleder:2002:FAB

[119] Mauro Steigleder. Factorization of the ashikhmin BRDF for real-time rendering. Journal of Graphics Tools: JGT, 7(4):61–68, 2002. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/Steigleder02/.

Wang:2002:FNM

[120] Yigang Wang, Bernd Fröhlich, and Martin Göbel. Fast normal map generation for simplified meshes. Journal of Graphics Tools: JGT, 7(4):69–82, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/WangFrohlichGobel02/.

Weiskopf:2002:RTD

[121] Daniel Weiskopf and Thomas Ertl. Realtime depth-cueing beyond fogging. Journal of Graphics Tools: JGT, 7(4):83–90, 2002. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/WeiskopfErtl02/.

Yang:2002:FIS

[122] Ruigang Yang and Greg Welch. Fast image segmentation and smoothing using commodity graphics hardware. Journal of Graphics Tools: JGT, 7(4):91–100, 2002. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/YangWelch02/.

Low:2003:CVF

[123] Kok-Lim Low and Adrian Ilie. Computing a view frustum to maximize an object’s image area. Journal of Graphics Tools: JGT, 8(1):3–15, 2003. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/LowIlie03/.

Shen:2003:FTT

[124] Hao Shen, Pheng Ann Heng, and Zesheng Tang. A fast triangle-triangle overlap test using signed distances. Journal of Graphics Tools: JGT, 8(1):16–24 (??), 2003. CODEN JGTOFD. ISSN

                                        1086-7651. URL http://www.acm.org/
                                        jgt/papers/ShenHengTang03/.

Guigue:2003:FRT

[125] Philippe Guigue and Olivier Devillers. Fast and robust triangle-triangle overlap test using orientation predicates. Journal of Graphics Tools: JGT, 8(1):25–32 (??), 2003. CODEN JGTOFD. ISSN

                                        1086-7651. URL http://www.acm.org/
                                        jgt/papers/GuigueDevillers03/.

Dawson:2003:CW

[126] Robert J. MacG. Dawson. Crooked wallpaper. Journal of Graphics Tools: JGT, 8(1):33–46, 2003. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

                                        acm.org/jgt/papers/Dawson03/.

Lewiner:2003:EIM

[127] Thomas Lewiner, Hélio Lopes, Antônio Wilson Vieira, and Geovan Tavares. Efficient implementation of marching cubes’ cases with topological guarantees. Journal of Graphics Tools: JGT, 8(2):1–15, 2003. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/LewinerEtAl03/.

Ward:2003:FRI

[128] Greg Ward. Fast, robust image registration for compositing high dynamic range photographcs from hand-held exposures. Journal of Graphics Tools: JGT, 8(2): 17–30, 2003. CODEN JGTOFD. ISSN

1086-7651. URL http://www.acm.org/
jgt/papers/Ward03/.

Kallay:2003:ITP

[129] Michael Kallay and Jason Lawrence. Improving the two-pass resampling algorithm. Journal of Graphics Tools: JGT, 8(2):31–40, 2003. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/KallayLawrence03/

.

Jianhui:2003:RBS

[130] Ye Jianhui. Region-based simplification of laser scanned meshes. Journal of Graphics Tools: JGT, 8(2):41–50, 2003. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/Jianhui03/.

Xue:2003:ESU

[131] Daqing Xue and Roger Crawfis. Efficient splatting using modern graphics hardware. Journal of Graphics Tools: JGT, 8(3):1–21, 2003. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

acm.org/jgt/papers/XueCrawfis03/.

Arvo:2003:OSM

[132] Jukka Arvo and Timo Aila. Optimized shadow mapping using the stencil buffer. Journal of Graphics Tools: JGT, 8(3): 23–32, 2003. CODEN JGTOFD. ISSN

                                        1086-7651. URL http://www.acm.org/
                                        jgt/papers/ArvoAila03/.

Adabala:2003:PTT

[133] Neeharika Adabala and Nadia Magnenat-Thalmann. A procedural thread texture model. Journal of Graphics Tools: JGT, 8(3):33–40, 2003. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/AdabalaMagnenatThalmann03/

.

Steigleder:2003:GSS

[134] Mauro Steigleder and Michael Mc-Cool. Generalized stratified sampling using the Hilbert curve. Journal of Graphics Tools: JGT, 8(3):41–47, 2003. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/SteiglederMcCool03/.

Mech:2003:RTI

[135] Radomír Mech. Real-time image-based rendering using surface proxies and texture packing. Journal of Graphics Tools: JGT, 8(4):1–19, 2003. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.acm.org/jgt/papers/Mech03/.

Akenine-Moller:2003:DVS

[136] Tomas Akenine-Möller and Ulf Assarsson. On the degree of vertices in a shadow volume silhouette. Journal of Graphics Tools: JGT, 8(4):21–24, 2003. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/AkenineMollerAssarsson03/.

Dur:2003:PLP

[137] Arne Dür and Sylvia Leimgruber.

A practical list-priority algorithm for 3D polygons. Journal of Graphics Tools: JGT, 8(4):25–36, 2003. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/DuerLeimgruber03/.

Platis:2003:FRT

[138] Nikos Platis and Theoharis Theoharis.

Fast ray-tetrahedron intersection using Plücker coordinates. Journal of Graphics Tools: JGT, 8(4):37–48, 2003. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/PlatisTheoharis03/.

McGuire:2004:OSS

[139] Morgan McGuire. Observations on silhouette sizes. Journal of Graphics Tools: JGT, 9(1):1–12, 2004. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/McGuire04/.

Wang:2004:SST

[140] Chung-Ming Wang and Nen-Chin Hwang. A stratified sampling technique for an ellipse. Journal of Graphics Tools: JGT, 9(1):13–22, 2004. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/WangHwang04/.

Telea:2004:IIT

[141] Alexandru Telea. An image inpainting technique based on the fast marching method. Journal of Graphics Tools: JGT, 9(1):23–34, 2004. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.acm.org/jgt/papers/Telea04/

.

Mahovsky:2004:FRA

[142] Jeffrey Mahovsky and Brian Wyvill. Fast ray-axis aligned bounding box overlap tests with Plücker coordinates. Journal of Graphics Tools: JGT, 9(1):35–46, 2004. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/MahovskyWyvill04/.

Schmidl:2004:CFU

[143] Harald Schmidl, Nolan Walker, and Ming Lin. CAB: Fast update of OBB trees for collision detection between articulated bodies. Journal of Graphics Tools: JGT, 9(2):1–9, 2004. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/SchmidlWalkerLin04/.

Barbier:2004:FDC

[144] Aurélien Barbier and Eric Galin. Fast distance computation between a point and cylinders, cones, line-swept spheres and cone-spheres. Journal of Graphics Tools: JGT, 9(2):11–19, 2004. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/BarbierGalin04/.

Velho:2004:DAM

[145] Luiz Velho. A dynamic adaptive mesh library based on stellar operators. Journal of Graphics Tools: JGT, 9(2):21–47, 2004. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://www.acm.org/jgt/
                                        papers/Velho04/.

Krishnaswamy:2004:IRC

[146] Aravind Krishnaswamy, Gladimir V. G. Baranoski, and Jon G. Rokne. Improving the reliability/cost ratio of goniophotometric comparisons. Journal of Graphics Tools: JGT, 9(3):1–20, 2004. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/BaranoskiEtAl04/.

Wang:2004:RFC

[147] Niniane Wang. Realistic and fast cloud rendering. Journal of Graphics Tools: JGT, 9(3):21–40, 2004. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.acm.org/jgt/papers/Wang04/.

Ramsey:2004:RBP

[148] Shaun D. Ramsey, Kristin Potter, and Charles Hansen. Ray bilinear patch intersections. Journal of Graphics Tools: JGT, 9(3):41–47, 2004. CODEN JGTOFD. ISSN 1086-

7651. URL http://www.acm.org/jgt/
papers/RamseyPotterHansen04/.

Stewart:2004:ADS

[149] Ian F. Stewart and André R. Foisy.

Arbitrary-degree subdivision with creases and attributes. Journal of Graphics Tools: JGT, 9(4):3–17, 2004. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/StewartFoisy04/.

Brentin:2004:MAS

[150] Greg Brentin and Greg Heflin. Maya^{T M} assets: A shot content management system. Journal of Graphics Tools: JGT, 9(4):19–31, 2004. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/BrentinHeflin04/.

Yee:2004:PMP

[151] Hector Yee. A perceptual metric for production testing. Journal of Graphics Tools: JGT, 9(4):33–40, 2004. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //jgt.akpeters.com/papers/Yee04/.

Piponi:2004:ADC

[152] Dan Piponi. Automatic differentiation, C++ templates, and photogrammetry. Journal of Graphics Tools: JGT, 9(4):41–55, 2004. CODEN JGTOFD. ISSN 1086-7651. URL http://jgt.

                                        akpeters.com/papers/Piponi04/.

Andersen:2004:CTP

[153] Ben Andersen and Kevin Noone. Controlled texture pushing and crossing seams in UV space using Maya and Photorealistic Renderman. Journal of Graphics Tools: JGT, 9(4):57–67, 2004. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://jgt.akpeters.
                                        com/papers/AndersenNoone04/.

Repenning:2005:IID

[154] Alexander Repenning. Inflatable icons: Diffusion-based interactive extrusion of 2D images into 3D models. Journal of Graphics Tools: JGT, 10(1):1–15, 2005. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://jgt.akpeters.
                                        com/papers/Repenning05/.

Lin:2005:PTD

[155] Zhouchen Lin, Hai-Tao Chen, Heung-Yeung Shum, and Jian Wang. Prefiltering two-dimensional polygons without clipping. Journal of Graphics Tools: JGT, 10(1):17–26, 2005. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://jgt.akpeters.
                                        com/papers/LinEtAl05/.

Lin:2005:OPF

[156] Zhouchen Lin, Hai-Tao Chen, Heung-Yeung Shum, and Jian Wang. Optimal polynomial filters. Journal of Graphics Tools: JGT, 10(1):27–38, 2005. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/LinEtAl05b/.

Stark:2005:GSS

[157] Michael Stark, Peter Shirley, and Michael Ashikhmin. Generation of stratified samples for B-spline pixel filtering. Journal of Graphics Tools: JGT, 10(1):39–48, 2005. CODEN JGTOFD. ISSN 1086-7651. URL http://jgt.

akpeters.com/papers/StarkEtAl05/.

Williams:2005:ERR

[158] Amy Williams, Steve Barrus, R. Keith Morley, and Peter Shirley. An efficient and robust ray-box intersection algorithm. Journal of Graphics Tools: JGT, 10(1):49–54, 2005. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/WilliamsEtAl05/.

Bethune:2005:ASG

[159] Christopher Bethune and A. James Stewart. Adaptive slice geometry for hardware-assisted volume rendering. Journal of Graphics Tools: JGT, 10(1):55–70, 2005. CODEN JGTOFD. ISSN 1086-7651. URL

http://jgt.akpeters.com/papers/
BethuneStewart05/.

Baerentzen:2005:HAP

[160] J. Andreas Bærentzen. Hardwareaccelerated point generation and rendering of point-based impostors. Journal of Graphics Tools: JGT, 10(2):1–12, 2005. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/Baerentzen05/.

Lofsted:2005:EFR

[161] Marta Löfsted and Tomas Akenine-Möller. An evaluation framework for ray–triangle intersection algorithms. Journal of Graphics Tools: JGT, 10(2):13–26, 2005. CODEN JGTOFD. ISSN 1086-7651. URL

                                        http://jgt.akpeters.com/papers/
                                        LofstedtAkenineMoller05/.

Malsch:2005:STD

[162] Elisabeth Anna Malsch, John Jeffy Lin, and Gautam Dasgupta. Smooth twodimensional interpolations: A recipe for all polygons. Journal of Graphics Tools: JGT, 10(2):27–39, 2005. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://jgt.akpeters.
                                        com/papers/MalschEtAl05/.

Reuter:2005:IAD

[163] Patrick Reuter, Johannes Behr, and Marc Alexa. An improved adjacency data structure for fast triangle stripping. Journal of Graphics Tools: JGT, 10(2):41–50, 2005. CODEN JGTOFD. ISSN 1086-7651. URL http://jgt.

                                        akpeters.com/papers/ReuterEtAl05/

.

Barrett:2005:EPE

[164] Sean Barrett. Efficient polygon edge enumeration. Journal of Graphics Tools: JGT, 10(2):51–53, 2005. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://jgt.akpeters.
                                        com/papers/Barrett05/.

Nijasure:2005:RTG

[165] Mangesh Nijasure, Sumanta N. Pattanaik, and Vineet Goel. Real-time global illumination on GPUs. Journal of Graphics Tools: JGT, 10(2):55–71, 2005. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/NijasureEtAl05/.

Akenine-Moller:2005:CTR

[166] Tomas Akenine-Möller and Timo Aila.

Conservative and tiled rasterization using a modified triangle set-up. Journal of Graphics Tools: JGT, 10(3):1–8, 2005. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/AkenineMollerAila05/.

Heidrich:2005:CBC

[167] Wolfgang Heidrich. Computing the barycentric coordinates of a projected point. Journal of Graphics Tools: JGT, 10(3):9–12, 2005. CODEN JGTOFD. ISSN 1086-7651. URL http://jgt.

akpeters.com/papers/Heidrich05/.

Chirkov:2005:FLS

[168] Nick Chirkov. Fast 3D line segment– triangle intersection test. Journal of Graphics Tools: JGT, 10(3):13–18, 2005. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/Chirkov05/.

Rokita:2005:DBS

[169] Przemyslaw Rokita. Depth-based selective antialiasing. Journal of Graphics Tools: JGT, 10(3):19–26, 2005. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/Rokita05/.

Segura:2005:EPC

[170] R. Segura, F. R. Feito, J. Ruiz de Miras, J. C. Torres, and C. Ogáyar. An efficient point classification algorithm for triangle meshes. Journal of Graphics Tools: JGT, 10(3):27–35, 2005. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://jgt.akpeters.
                                        com/papers/SeguraEtAl05/.

Buss:2005:SDL

[171] Samuel R. Buss and Jin-Su Kim. Selectively damped least squares for inverse kinematics. Journal of Graphics Tools: JGT, 10(3):37–49, 2005. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://jgt.akpeters.
                                        com/papers/BussKim05/.

Singh:2005:IIC

[172] Karan Singh and Cindy Grimm. Implementing the IBar camera widget. Journal of Graphics Tools: JGT, 10 (3):51–64, 2005. CODEN JGTOFD. ISSN 1086-7651. URL http://jgt.

                                        akpeters.com/papers/SinghGrimm05/

.

Toksvig:2005:MNM

[173] Michael Toksvig. Mipmapping normal maps. Journal of Graphics Tools: JGT, 10(3):65–71, 2005. CODEN JGTOFD. ISSN 1086-7651. URL http://jgt.

                                        akpeters.com/papers/Toksvig05/.

Wrotek:2005:RTC

[174] Pawel Wrotek, Alexander Rice, and Morgan McGuire. Real-time collision deformations using graphics hardware. Journal of Graphics Tools: JGT, 10 (4):1–22, 2005. CODEN JGTOFD. ISSN 1086-7651. URL http://jgt.

                                        akpeters.com/papers/WrotekEtAl05/

.

Lagae:2005:ERQ

[175] Ares Lagae and Philip Dutré. An efficient ray-quadrilateral intersection test. Journal of Graphics Tools: JGT, 10 (4):23–32, 2005. CODEN JGTOFD. ISSN 1086-7651. URL http://jgt.

akpeters.com/papers/LagaeDutre05/

.

Dingliana:2005:VBA

[176] John Dingliana and Carol O’Sullivan.

A voxel-based approach to approximate collision handling. Journal of Graphics Tools: JGT, 10(4):33–48, 2005. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/DinglianaOSullivan05/.

Patel:2005:SDF

[177] Mayur Patel and Noah Taylor. Simple divergence-free fields for artistic simulation. Journal of Graphics Tools: JGT, 10(4):49–60, 2005. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/PatelTaylor05/.

Lum:2005:CHR

[178] Eric B. Lum, Kwan-Liu Ma, and Nelson Max. Calculating hierarchical radiosity form factors using programmable graphics hardware. Journal of Graphics Tools: JGT, 10(4):61–71, 2005. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/LumEtAl05/.

Lacewell:2006:SBC

[179] J. Dylan Lacewell, Dave Edwards, Peter Shirley, and William B. Thompson. Stochastic billboard clouds for interactive foliage rendering. Journal of Graphics Tools: JGT, 11(1):1–12, 2006. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/LacewellEtAl06/.

Ben-Artzi:2006:ESS

[180] Aner Ben-Artzi, Ravi Ramamoorthi, and Maneesh Agrawala. Efficient shadows from sampled environment maps. Journal of Graphics Tools: JGT, 11(1):13–36, 2006. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://jgt.akpeters.
                                        com/papers/BenArtziEtAl06/.

Ma:2006:HPM

[181] YingLiang Ma, W. T. Hewitt, and Martin Turner. A high-performance method for calculating the minimum distance between two 2D and 3D NURBS curves. Journal of Graphics Tools: JGT, 11(1):37–50, 2006. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://jgt.akpeters.
                                        com/papers/MaHewittTurner06/.

Dur:2006:INW

[182] Arne Dür. An improved normalization for the Ward reflectance model. Journal of Graphics Tools: JGT, 11 (1):51–59, 2006. CODEN JGTOFD. ISSN 1086-7651. URL http://jgt.

                                        akpeters.com/papers/Dur06/.

Tumblin:2006:EDI

[183] Jack Tumblin. Exact 2-D integration inside quadrilateral boundaries. Journal of Graphics Tools: JGT, 11(1):61–71, 2006. CODEN JGTOFD. ISSN 1086-

                                        7651. URL http://jgt.akpeters.
                                        com/papers/Tumblin06/.

Benthin:2006:TIR

[184] Carsten Benthin, Ingo Wald, and Philipp Slusallek. Techniques for interactive ray tracing of Bézier surfaces. Journal of Graphics Tools:

JGT, 11(2):1–16, 2006. CODEN JGTOFD. ISSN 1086-7651. URL

http://jgt.akpeters.com/papers/
BenthinEtAl06/.

Mousa:2006:DSH

[185] M. Mousa, R. Chaine, and S. Akkouche.

Direct spherical harmonic transform of a triangulated mesh. Journal of Graphics Tools: JGT, 11(2):17–26, 2006. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/MousaEtAl06/.

Jones:2006:EGP

[186] Thouis R. Jones. Efficient generation of Poisson-disk sampling patterns. Journal of Graphics Tools: JGT, 11(2):27–36, 2006. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/Jones05/.

Redon:2006:FML

[187] Stephane Redon and Ming C. Lin.

A fast method for local penetration depth computation. Journal of Graphics Tools: JGT, 11(2):37–50, 2006. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/RedonLin06/.

Kallay:2006:CMI

[188] Michael Kallay. Computing the moment of inertia of a solid defined by a triangle mesh. Journal of Graphics Tools: JGT, 11(2):51–57, 2006. CODEN JGTOFD. ISSN 1086-7651. URL http://jgt.

akpeters.com/papers/Kallay06/.

deBoer:2006:SPT

[189] Willem H. de Boer. Smooth penumbra transitions with shadow maps. Journal of Graphics Tools: JGT, 11(2):59–71, 2006. CODEN JGTOFD. ISSN 1086-

7651. URL http://jgt.akpeters.
com/papers/deBoer06/.

Kontkanen:2006:SPV

[190] Janne Kontkanen and Samuli Laine. Sampling precomputed volumetric lighting. Journal of Graphics Tools: JGT, 11(3):1–16, ???? 2006. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=11&issue=3&spage=1.

Li:2006:BGA

[191] Zhong Li and Lizhuang Ma. A bidirectional generating algorithm for rational parametric curves. Journal of Graphics Tools: JGT, 11 (3):17–26, ???? 2006. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=11&issue=3&spage=17.

Zuniga:2006:RQW

[192] Miguel R. Zuniga and Jeffrey K. Uhlmann. Ray queries with wide object isolation and the DE-tree. Journal of Graphics Tools: JGT, 11 (3):27–45, ???? 2006. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=11&issue=3&spage=27.

Sanderson:2006:ARD

[193] Allen R. Sanderson, Robert M. Kirby, Chris R. Johnson, and Lingfa Yang. Advanced reaction-diffusion models for texture synthesis. Journal of Graphics Tools: JGT, 11

(3):47–71, ????  2006.  CODEN

JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=11&issue=3&spage=47.

Bernardon:2006:GBT

[194] Fábio F. Bernardon, Christian A. Pagot, João L. D. Comba, and Cláudio T. Silva. GPU-based tiled ray casting using depth peeling. Journal of Graphics Tools: JGT, 11(4):1–16, ???? 2006. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=11&issue=4&spage=1.

Liu:2006:HRI

[195] Peiran Liu, Nicolas D. Georganas, and Gerhard Roth. Handling rapid interference detection of progressive meshes using active bounding trees. Journal of Graphics Tools: JGT, 11 (4):17–37, ???? 2006. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=11&issue=4&spage=17.

Fischer:2006:FAH

[196] Ian Fischer and Craig Gotsman. Fast approximation of high-order Voronoi diagrams and distance transforms on the GPU. Journal of Graphics Tools: JGT, 11(4):39–60, ???? 2006. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=11&issue=4&spage=39.

Cline:2006:LBV

[197] David Cline, Kevin Steele, and Parris Egbert. Lightweight bounding volumes for ray tracing. Journal of Graphics Tools: JGT, 11 (4):61–71, ???? 2006. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=11&issue=4&spage=61.

Larsson:2007:FSB

[198] Thomas Larsson, Tomas Akenine-Möller, and Eric Lengyel. On faster sphere-box overlap testing. Journal of Graphics Tools: JGT, 12 (1):3–8, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=1&spage=3.

Stephenson:2007:IMB

[199] Ian Stephenson. Improving motion blur: Shutter efficiency and temporal sampling. Journal of Graphics Tools: JGT, 12(1):9–15, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=1&spage=9.

Friedel:2007:USP

[200] Ilja Friedel, Peter Schröder, and Mathieu Desbrun. Unconstrained spherical parameterization. Journal of Graphics Tools: JGT, 12 (1):17–26, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=1&spage=17.

Marinov:2007:GBM

[201] Martin Marinov, Mario Botsch, and Leif Kobbelt. GPU-based multiresolution deformation using approximate normal field reconstruction. Journal of Graphics Tools: JGT, 12 (1):27–46, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=1&spage=27.

Arvo:2007:AFS

[202] Jukka Arvo. Alias-free shadow maps using graphics hardware. Journal of Graphics Tools: JGT, 12 (1):47–59, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=1&spage=47.

Passalis:2007:GVA

[203] Georgios Passalis, Theoharis Theoharis, George Toderici, and Ioannis A. Kakadiaris. General voxelization algorithm with scalable GPU implementation. Journal of Graphics Tools: JGT, 12(1):61–71, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=1&spage=61.

Arvo:2007:SSC

[204] James Arvo and Kevin Novins. Stratified sampling of convex quadrilaterals. Journal of Graphics Tools: JGT, 12(2):1–12, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=2&spage=1.

Bradley:2007:ATU

[205] Derek Bradley and Gerhard Roth. Adaptive thresholding using the integral image. Journal of Graphics Tools: JGT, 12(2):13–21, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=12&issue=2&spage=13.

Lengyel:2007:UDF

[206] Eric Lengyel. Unified distance formulas for halfspace fog. Journal of Graphics Tools: JGT, 12 (2):23–32, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=12&issue=2&spage=23.

Max:2007:HPC

[207] Nelson Max. Hexahedron projection for curvilinear grids. Journal of Graphics Tools: JGT, 12 (2):33–45, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=12&issue=2&spage=33.

Karasuda:2007:DRT

[208] Eric Karasuda and Sara McMains. Displaying readable text in a headtracked, stereoscopic virtual environment. Journal of Graphics Tools: JGT, 12(2):47–57, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=12&issue=2&spage=47.

Malmer:2007:FPA

[209] Mattias Malmer, Fredrik Malmer, Ulf Assarsson, and Nicolas Holzschuch. Fast precomputed ambient occlusion for proximity shadows. Journal of Graphics Tools: JGT, 12 (2):59–71, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=2&spage=59.

Li:2007:IIA

[210] Xin Li. iSlerp: An incremental approach to Slerp. Journal of Graphics Tools: JGT, 12(3):1–6, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=3&spage=1.

Lacewell:2007:EEC

[211] Dylan Lacewell and Brent Burley. Exact evaluation of Catmull–Clark subdivision surfaces near B-spline boundaries. Journal of Graphics Tools: JGT, 12(3):7–15, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=3&spage=7.

Risser:2007:FRM

[212] Eric Risser, Musawir Shah, and Sumanta Pattanaik. Faster relief mapping using the secant method. Journal of Graphics Tools: JGT, 12 (3):17–24, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=3&spage=17.

McDonnell:2007:PFP

[213] Kevin T. McDonnell and Hong Qin. PB-FFD: A point-based technique for free-form deformation. Journal of Graphics Tools: JGT, 12 (3):25–41, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=12&issue=3&spage=25.

Mikkelsen:2007:SPP

[214] Morten S. Mikkelsen. Separatingplane perspective shadow mapping. Journal of Graphics Tools: JGT, 12 (3):43–54, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=12&issue=3&spage=43.

Sussner:2007:BAS

[215] Gerd Sußner, Marc Stamminger, and Günther Greiner. Bidirectional adaptive \sqrt{3}-subdivision. Journal of Graphics Tools: JGT, 12 (4):1–24, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=12&issue=4&spage=1.

Grimm:2007:CCM

[216] Cindy M. Grimm and Bill Niebruegge. Continuous cube mapping. Journal of Graphics Tools: JGT, 12 (4):25–34, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=12&issue=4&spage=25.

Eisemann:2007:FRA

[217] Martin Eisemann, Marcus Magnor, Thorsten Grosch, and Stefan Müller. Fast ray/axis-aligned bounding box overlap tests using ray slopes. Journal of Graphics Tools: JGT, 12 (4):35–46, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=4&spage=35.

Parus:2007:FCV

[218] Jindrich Parus, Ivana Kolingerová, and Anders Hast. Fast computation of vertex normals for linearly deforming meshes. Journal of Graphics Tools: JGT, 12(4):47–58, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=4&spage=47.

Hasselgren:2007:TSV

[219] Jon Hasselgren and Tomas Akenine-Möller. Textured shadow volumes. Journal of Graphics Tools: JGT, 12 (4):59–72, ???? 2007. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=12&issue=4&spage=59.

Wyman:2008:RNI

[220] Chris Wyman and Carsten Dachsbacher. Reducing noise in imagespace caustics with variable-sized splatting. Journal of Graphics Tools: JGT, 13(1):1–17, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=1&spage=1.

Bavoil:2008:RSS

[221] Louis Bavoil, Steven P. Callahan, and Cláudio T. Silva. Robust soft shadow mapping with backprojection and depth peeling. Journal of Graphics Tools: JGT, 13(1):19–30, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=13&issue=1&spage=19.

Larsson:2008:EEO

[222] Thomas Larsson. An efficient ellipsoid-OBB intersection test. Journal of Graphics Tools: JGT, 13(1): 31–43, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=13&issue=1&spage=31.

DeCoro:2008:SSF

[223] Christopher DeCoro and Szymon Rusinkiewicz. Subtractive shadows: A flexible framework for shadow level of detail. Journal of Graphics Tools: JGT, 13(1):45–56, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=13&issue=1&spage=45.

Hutchinson:2008:PCT

[224] Elena Jakubiak Hutchinson, Sarah Frisken, and Ronald Perry. Proximity cluster trees. Journal of Graphics Tools: JGT, 13(1):57–69, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=13&issue=1&spage=57.

Hausner:2008:VDH

[225] Alejo Hausner. Versatile decorative halftoning. Journal of Graphics Tools: JGT, 13(2):1–12, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=2&spage=1.

Faust:2008:ECG

[226] Martin Faust. Enhanced color-to-gray conversion. Journal of Graphics Tools: JGT, 13(2):13–19, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=2&spage=13.

Vidal:2008:SES

[227] Vincent Vidal, Xing Mei, and Philippe Decaudin. Simple empty-space removal for interactive volume rendering. Journal of Graphics Tools: JGT, 13(2):21–36, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=2&spage=21.

Frisken:2008:ECF

[228] Sarah F. Frisken. Efficient curve fitting. Journal of Graphics Tools: JGT, 13(2):37–54, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=2&spage=37.

Wang:2008:LCM

[229] Wencheng Wang, Chunjuan Sun, Jing Li, and Enhua Wu. Line clipping by managing polygon edges in convex polylines. Journal of Graphics Tools: JGT, 13(2):55–71, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=2&spage=55.

Kim:2008:SVA

[230] Byungmoon Kim, Kihwan Kim, and Greg Turk. A shadow-volume algorithm for opaque and transparent nonmanifold casters. Journal of Graphics Tools: JGT, 13(3):1–14, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=3&spage=1.

Silva:2008:NNB

[231] Frutuoso G. M. Silva and Abel J. P. Gomes. A nonoriented, nonmanifold boundary representation for geometric applications. Journal of Graphics Tools: JGT, 13(3):15–33, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=3&spage=15.

Brodu:2008:QSI

[232] Nicolas Brodu. Query sphere indexing for neighborhood requests. Journal of Graphics Tools: JGT, 13 (3):35–51, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=3&spage=35.

Clarberg:2008:FEA

[233] Petrik Clarberg. Fast equal-area mapping of the (hemi)sphere using SIMD. Journal of Graphics Tools: JGT, 13

(3):53–68, ????  2008.  CODEN

JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=3&spage=53.

McGuire:2008:EHQ

[234] Morgan McGuire. Efficient, highquality Bayer demosaic filtering on GPUs. Journal of Graphics Tools: JGT, 13(4):1–16, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=4&spage=1.

Rousseau:2008:GR

[235] Pierre Rousseau, Vincent Jolivet, and Djamchid Ghazanfarpour. GPU rainfall. Journal of Graphics Tools: JGT, 13(4):17–33, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=4&spage=17.

Ebner:2008:GCC

[236] Marc Ebner. GPU color constancy.

Journal of Graphics Tools: JGT, 13 (4):35–51, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=4&spage=35.

Bai:2008:CCM

[237] Linge Bai and David Breen. Calculating center of mass in an unbounded 2D environment. Journal of Graphics Tools: JGT, 13(4):53–60, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=2151-237X&
volume=13&issue=4&spage=53.

Ruijters:2008:EGB

[238] Daniel Ruijters, Bart M. ter Haar Romeny, and Paul Suetens. Efficient GPU-based texture interpolation using uniform Bsplines. Journal of Graphics Tools: JGT, 13(4):61–69, ???? 2008. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=13&issue=4&spage=61.

Callahan:2009:AUV

[239] Steven P. Callahan and Cláudio T. Silva. Accelerating unstructured volume rendering with joint bilateral upsampling. Journal of Graphics, GPU, and Game Tools, 14(1):1–15, ???? 2009. CO-DEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=1&spage=1.

Igarashi:2009:IRP

[240] Takeo Igarashi and Yuki Igarashi. Implementing as-rigid-as-possible shape manipulation and surface flattening. Journal of Graphics, GPU, and Game Tools, 14(1):17–30, ???? 2009. CO-DEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=1&spage=17.

Bayraktar:2009:GBN

[241] Serkan Bayraktar, Uğur Güdükbay, and Bülent Özgüç. GPU-based neighborsearch algorithm for particle simulations. Journal of Graphics, GPU, and Game Tools, 14(1):31–42, ???? 2009. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=1&spage=31.

Mantiuk:2009:VHD

[242] Rafal Mantiuk and Wolfgang Heidrich. Visualizing high dynamic range images in a Web browser. Journal of Graphics, GPU, and Game Tools, 14(1):43–53, ???? 2009. CO-DEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=1&spage=43.

Stark:2009:ECP

[243] Michael M. Stark. Efficient construction of perpendicular vectors without branching. Journal of Graphics, GPU, and Game Tools, 14 (1):55–62, ???? 2009. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=1&spage=55.

Piponi:2009:TTP

[244] Dan Piponi. Two tricks for the price of one: Linear filters and their transposes. Journal of Graphics, GPU, and Game Tools, 14(1):63–72, ???? 2009. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=1&spage=63.

Stark:2009:FSC

[245] Michael M. Stark. Fast and stable conformal mapping between a disc and a square. Journal of Graphics, GPU, and Game Tools, 14(2):1–23, ???? 2009. CO-DEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=14&issue=2&spage=1.

Dammertz:2009:PSL

[246] Holger Dammertz and Johannes Hanika. Plane sampling for light paths from the environment map. Journal of Graphics, GPU, and Game Tools, 14 (2):25–31, ???? 2009. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=14&issue=2&spage=25.

Sadeghi:2009:CPT

[247] Iman Sadeghi, Bin Chen, and Henrik Wann Jensen. Coherent path tracing. Journal of Graphics, GPU, and Game Tools, 14(2):33–43, ???? 2009. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=14&issue=2&spage=33.

Pouchol:2009:HHS

[248] Mickael Pouchol, Alexandre Ahmad, Benoit Crespin, and Olivier Terraz. A hierarchical hashing scheme for nearest neighbor search and broadphase collision detection. Journal of Graphics, GPU, and Game Tools, 14(2):45–59, ???? 2009. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=2&spage=45.

Augsdorfer:2009:RPR

[249] Ursula H. Augsdörfer, Neil A. Dodgson, and Malcolm A. Sabin. Removing polar rendering artifacts in subdivision surfaces. Journal of Graphics, GPU, and Game Tools, 14(2):61–76, ???? 2009. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=2&spage=61.

Cords:2009:ISS

[250] Hilko Cords and Oliver G. Staadt. Interactive screen-space surface rendering of dynamic particle clouds. Journal of Graphics, GPU, and Game Tools, 14(3):1–19, ???? 2009. CO-DEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=3&spage=1.

Forest:2009:RTH

[251] Vincent Forest, Loic Barthe, and Mathias Paulin. Real-time hierarchical binary-scene voxelization. Journal of Graphics, GPU, and Game Tools, 14(3):21–34, ???? 2009. CO-DEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=3&spage=21.

Schoning:2009:BIM

[252] Johannes Schöning, Jonathan Hook, Nima Motamedi, Patrick Olivier, Florian Echtler, Peter Brandl, Laurence Muller, Florian Daiber, Otmar Hilliges, Markus Loechtefeld, Tim Roth, Dominik Schmidt, and Ulrich von Zadow. Building interactive multi-touch surfaces. Journal of Graphics, GPU, and Game Tools, 14(3):35–55, ???? 2009. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=14&issue=3&spage=35.

Starck:2009:FVV

[253] J. Starck, J. Kilner, and A. Hilton. A free-viewpoint video renderer. Journal of Graphics, GPU, and Game Tools, 14(3):57–72, ???? 2009. CO-DEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=14&issue=3&spage=57.

vanWelbergen:2009:RTA

[254] Herwin van Welbergen, Job Zwiers, and Zsófia M. Ruttkay. Real-time animation using a mix of physical simulation and kinematics. Journal of Graphics, GPU, and Game Tools, 14(4):1–21, ???? 2009. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=2151-237X&
                                        volume=14&issue=4&spage=1.

McGuire:2009:ILS

[255] Morgan McGuire. An inexpensive light stage dome. Journal of Graphics, GPU, and Game Tools, 14 (4):23–29, ???? 2009. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=4&spage=23.

Heinly:2009:IRT

[256] Jared Heinly, Shawn Recker, Kevin Bensema, Jesse Porch, and Christiaan Gribble. Integer ray tracing. Journal of Graphics, GPU, and Game Tools, 14(4):31–56, ???? 2009. CO-DEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=4&spage=31.

Strugar:2009:CDD

[257] Filip Strugar. Continuous distancedependent level of detail for rendering heightmaps. Journal of Graphics, GPU, and Game Tools, 14 (4):57–74, ???? 2009. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=2151-237X&
volume=14&issue=4&spage=57.

Boulanger:2010:HFS

[258] Kevin Boulanger, Kadi Bouatouch, and Sumanta Pattanaik. High-frequency shadows for real-time rendering of trees. Journal of Graphics, GPU, and Game Tools, 15(1):1–12, 2010. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Brunton:2010:SHT

[259] Alan Brunton, Jochen Lang, and Eric Dubois. Spherical harmonic transforms and convolutions on the GPU. Journal of Graphics, GPU, and Game Tools, 15(1):13–27, 2010. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Ascencio-Lopez:2010:AIS

[260] Ignacio Ascencio-Lopez, Oscar Meruvia-Pastor, and Hugo Hidalgo-Silva. Adaptive incremental stippling using the Poisson-disk distribution. Journal of Graphics, GPU, and Game Tools, 15(1): 29–47, 2010. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Mikkelsen:2010:BMU

[261] Morten S. Mikkelsen. Bump mapping unparametrized surfaces on the GPU. Journal of Graphics, GPU, and Game Tools, 15(1):49–61, 2010. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Sherman:2010:TPT

[262] Glen Aldridge Sherman. A triangle product for texture stretch. Journal of Graphics, GPU, and Game Tools, 15(1): 63–72, 2010. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Ahmed:2011:PPQ

[263] Abdalla G. M. Ahmed. Pixel patterns from quantization artifacts of forward affine mapping. Journal of Graphics, GPU, and Game Tools, 15(2):73–94, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Glassner:2011:RCE

[264] Andrew Glassner. Reconciling circular and elliptical arcs. Journal of Graphics, GPU, and Game Tools, 15(2):95–98, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Lengyel:2011:TCD

[265] Eric Lengyel. Transition cells for dynamic multiresolution marching cubes. Journal of Graphics, GPU, and Game Tools, 15(2):99–122, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Munkberg:2011:BCM

[266] Jacob Munkberg and Tomas Akenine-Möller. Backface culling for motion blur and depth of field. Journal of Graphics, GPU, and Game Tools, 15(2):123–139, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Ernst:2011:RTS

[267] Manfred Ernst and Sven Woop. Ray tracing with shared-plane bounding volume hierarchies. Journal of Graphics, GPU, and Game Tools, 15(3):141–151, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Schlomer:2011:ASA

[268] Thomas Schlömer and Oliver Deussen.

Accurate spectral analysis of twodimensional point sets. Journal of Graphics, GPU, and Game Tools, 15(3): 152–160, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Eberly:2011:FAA

[269] David Eberly. A fast and accurate algorithm for computing SLERP. Journal of Graphics, GPU, and Game Tools, 15(3):161–176, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Jones:2011:LTP

[270] Thouis R. Jones and David R. Karger. Linear-time Poisson-disk patterns. Journal of Graphics, GPU, and Game Tools, 15(3):177–182, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Pegoraro:2011:ECV

[271] Vincent Pegoraro and Philipp Slusallek. On the evaluation of the complex-valued exponential integral. Journal of Graphics, GPU, and Game Tools, 15(3):183– 198, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Shirley:2011:SAM

[272] Peter Shirley, Austin Robison, and R. Keith Morley. A simple algorithm for managing color in global tone reproduction. Journal of Graphics, GPU, and Game Tools, 15(3):199–205, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

McGuire:2011:E

[273] Morgan McGuire. Editorial. Journal of Graphics, GPU, and Game Tools, 15(4):207–209, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Cowan:2011:GBM

[274] Brent Cowan and Bill Kapralos. A GPUbased method to approximate acoustical reflectivity. Journal of Graphics, GPU, and Game Tools, 15(4):210–215, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

McGuire:2011:ETQ

[275] Morgan McGuire. Efficient triangle and quadrilateral clipping within shaders. Journal of Graphics, GPU, and Game Tools, 15(4):216–224, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Thomsen:2011:ARU

[276] Andreas Thomsen and Kasper Høy Nielsen. Approximate radiosity using stochastic depth buffering. Journal of Graphics, GPU, and Game Tools, 15(4):225–234, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Olsson:2011:TS

[277] Ola Olsson and Ulf Assarsson. Tiled shading. Journal of Graphics, GPU, and Game Tools, 15(4):235–251, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Sloan:2011:WS

[278] Peter-Pike Sloan, Derek Nowrouzezahrai, and Hong Yuan. Wrap shading. Journal of Graphics, GPU, and Game Tools, 15 (4):252–259, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Anonymous:2011:EBE

[279] Anonymous. Editorial board EOV. Journal of Graphics, GPU, and Game Tools, 15(4):ebi–??, 2011. CODEN ???? ISSN 2151-2272 (print), 2151-237X (electronic).

Yuan:2012:IR

[280] Hong Yuan, Derek Nowrouzezahrai, and Peter-Pike Sloan. Irradiance rigs. Journal of Graphics Tools: JGT, 16(1):1–11, 2012. CODEN JGTOFD. ISSN 1086-7651.

Ray:2012:AAL

[281] Bimal Kumar Ray. An alternative algorithm for line clipping. Journal of Graphics Tools: JGT, 16(1):12–24, 2012. CODEN JGTOFD. ISSN 1086-7651.

Vo:2012:SEM

[282] Huy T. Vo, Claudio T. Silva, Luiz F. Scheidegger, and Valerio Pascucci. Simple and efficient mesh layout with spacefilling curves. Journal of Graphics Tools: JGT, 16(1):25–39, 2012. CODEN JGTOFD. ISSN 1086-7651.

Upchurch:2012:TPP

[283] Paul Upchurch and Mathieu Desbrun. Tightening the precision of perspective rendering. Journal of Graphics Tools: JGT, 16(1):40–56, 2012. CODEN JGTOFD. ISSN 1086-7651.

LeBlanc:2012:MBF

[284] Graham LeBlanc, Andrew Shouldice, Dirk V. Arnold, and Stephen Brooks. Multi-band Fourier synthesis of ocean waves. Journal of Graphics Tools: JGT, 16(2):57–70, 2012. CODEN JGTOFD. ISSN 1086-7651.

Iwanicki:2012:NML

[285] Michal Iwanicki and Peter-Pike Sloan.

Normal mapping with low-frequency precomputed visibility. Journal of Graphics Tools: JGT, 16(2):71–84, 2012. CODEN JGTOFD. ISSN 1086-7651.

McEwan:2012:ECN

[286] Ian McEwan, David Sheets, Mark Richardson, and Stefan Gustavson. Efficient computational noise in GLSL. Journal of Graphics Tools: JGT, 16(2): 85–94, 2012. CODEN JGTOFD. ISSN 1086-7651.

Schretter:2012:GRS

[287] Colas Schretter, Leif Kobbelt, and Paul-Olivier Dehaye. Golden ratio sequences for low-discrepancy sampling. Journal of Graphics Tools: JGT, 16(2):95–104, 2012. CODEN JGTOFD. ISSN 1086-7651.

Pfeifle:2012:RCC

[288] Ron Pfeifle. Rendering cubic curves on a GPU with Floater’s implicitization. Journal of Graphics Tools: JGT, 16 (2):105–122, 2012. CODEN JGTOFD. ISSN 1086-7651.

Nunes:2012:RTD

[289] Gustavo Nunes, Alexandre Valdetaro, Alberto Raposo, Bruno Feijó, and Rodrigo de Toledo. Rendering tubes from discrete curves using hardware tessellation. Journal of Graphics Tools: JGT, 16(3):123–143, 2012. CODEN JGTOFD. ISSN 1086-7651.

Tarini:2012:CTP

[290] Marco Tarini. Cylindrical and toroidal parameterizations without vertex seams.

Journal of Graphics Tools: JGT, 16 (3):144–150, 2012. CODEN JGTOFD. ISSN 1086-7651.

Frisvad:2012:BOB

[291] Jeppe Revall Frisvad. Building an orthonormal basis from a 3D unit vector without normalization. Journal of Graphics Tools: JGT, 16(3):151–159, 2012. CODEN JGTOFD. ISSN 1086-7651.

Anjyo:2012:PAD

[292] Ken Anjyo, Hideki Todo, and J. P. Lewis. A practical approach to direct manipulation blendshapes. Journal of Graphics Tools: JGT, 16(3):160–176, 2012. CODEN JGTOFD. ISSN 1086-7651.

Kenwright:2012:IKC

[293] Ben Kenwright. Inverse kinematics — cyclic coordinate descent (CCD). Journal of Graphics Tools: JGT, 16(4):177– 217, 2012.

Champagnat:2012:ECB

[294] Frédéric Champagnat and Yves Le Sant. Efficient cubic B-spline image interpolation on a GPU. Journal of Graphics Tools: JGT, 16(4):218–232, 2012. See erratum [302].

Seo:2012:LST

[295] Jaewoo Seo and Ken Anjyo. Line selection tool for 3D artists. Journal of Graphics Tools: JGT, 16(4):233–244, 2012.

Anonymous:2012:EBE

[296] Anonymous. Editorial board EOV. Journal of Graphics Tools: JGT, 16(4): ebi, 2012.

Jacobson:2013:BMG

[297] Alec Jacobson. Bijective mappings with generalized barycentric coordinates: A counterexample. Journal of Graphics Tools: JGT, 17(1–2):1–4, 2013. CODEN JGTOFD. ISSN 1086-7651.

Manoharan:2013:ICD

[298] Prabukumar Manoharan and Bimal Kumar Ray. An improved circle drawing algorithm on a hexagonal grid. Journal of Graphics Tools: JGT, 17(1–2):5–15, 2013. CODEN JGTOFD. ISSN 1086-7651.

Frogley:2013:FRD

[299] D. Frogley and M. D. Jones. Fast relabeling of deformable Delaunay tetrahedral meshes using a compact uniform grid. Journal of Graphics Tools: JGT, 17(1–2):17–29, 2013. CODEN JGTOFD. ISSN 1086-7651.

Mousas:2013:MEE

[300] Christos Mousas, Paul Newbury, and Christos-Nikolaos Anagnostopoulos. The minimum energy expenditure shortest path method. Journal of Graphics Tools: JGT, 17(1–2):31–44, 2013. CO-DEN JGTOFD. ISSN 1086-7651.

Pobegailo:2013:CSC

[301] Alexander P. Pobegailo. Construction of small circular arcs on a sphere of unit quaternions. Journal of Graphics Tools: JGT, 17(1–2):45–51, 2013. CO-DEN JGTOFD. ISSN 1086-7651.

Anonymous:2013:EEC

[302] Anonymous. Erratum: “Efficient Cubic B-spline Image Interpolation on a GPU”, by Frédéric Champagnat and Yves Le Sant, J. Graph. Tools 14(4)

218–232 (2012). Journal of Graphics Tools: JGT, 17(1–2):53, 2013. CODEN JGTOFD. ISSN 1086-7651. See [294].

Obaid:2013:MES

[303] Mohammad Obaid, Erik Sintorn, Daniel Sjölie, and Morten Fjeld. Message from the Editors: SIGRAD 2014 special issue of the Journal of Graphics Tools. Journal of Graphics Tools: JGT, 17 (3):55–58, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

tandfonline.com/doi/abs/10.1080/
2165347X.2015.1043848.

Fratarcangeli:2013:GBI

[304] Marco Fratarcangeli and Fabio Pellacini. A GPU-based implementation of position based dynamics for interactive deformable bodies. Journal of Graphics Tools: JGT, 17(3):59–66, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.tandfonline.com/doi/abs/10.
1080/2165347X.2015.1030525.

Kallberg:2013:FAM

[305] Linus Källberg and Thomas Larsson. Faster approximation of minimum enclosing balls by distance filtering and GPU parallelization. Journal of Graphics Tools: JGT, 17(3): 67–84, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

tandfonline.com/doi/abs/10.1080/
2165347X.2015.1037471.

Marreiros:2013:SMM

[306] Filipe M. M. Marreiros and Örjan Smedby. Single-monitor-mirror stereoscopic display. Journal of Graphics Tools: JGT, 17(3):85–97, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.tandfonline.com/doi/abs/10.
1080/2165347X.2015.1028690.

Dupre:2013:PPS

[307] R. Dupre, V. Argyriou, and D. Greenhill. Prediction of physics simulation using dimensionality reduction and regression. Journal of Graphics Tools: JGT, 17(3):99–110, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.tandfonline.com/doi/abs/10.
1080/2165347X.2015.1034813.

Banterle:2013:E

[308] Francesco Banterle. Editorial. Journal of Graphics Tools: JGT, 17 (4):111, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

tandfonline.com/doi/abs/10.1080/
2165347X.2015.1046323.

Asuni:2013:TLD

[309] Nicola Asuni and Andrea Giachetti.

TESTIMAGES: A large data archive for display and algorithm testing. Journal of Graphics Tools: JGT, 17(4): 113–125, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

tandfonline.com/doi/abs/10.1080/
2165347X.2015.1024298.

Keinert:2013:IRC

[310] Benjamin Keinert, Henry Schäfer, Johann Korndörfer, Urs Ganse, and Marc Stamminger. Improved ray casting of procedural distance bounds. Journal of Graphics Tools: JGT, 17(4): 127–138, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

tandfonline.com/doi/abs/10.1080/
2165347X.2015.1033069.

Ivo:2013:ASB

[311] Rafael Ivo, Fabio Ganovelli, Creto Vidal, Joaquim Bento Cavalcante-Neto, and Roberto Scopigno. Adapting splatbased models to curved sharp features. Journal of Graphics Tools: JGT, 17 (4):139–150, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        2165347X.2015.1034331.

Schmidt:2013:TIB

[312] Michael R. Schmidt. Toward improved batchability of 3D objects using a consolidated shader. Journal of Graphics Tools: JGT, 17(4): 151–158, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        2165347X.2014.909340.

Schulz:2013:CAT

[313] Adriana Schulz, Wojciech Matusik, and Luiz Velho. ChoreoGraphics: An authoring tool for dance shows. Journal of Graphics Tools: JGT, 17(4): 159–176, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        2165347X.2014.909341.

Anonymous:2013:EEB

[314] Anonymous. EOV editorial board. Journal of Graphics Tools: JGT, 17 (4):ebi, 2013. CODEN JGTOFD. ISSN 1086-7651. URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        2165347X.2013.1064679.